A Complete Bibliography of Publications in the Journal of Digital Information

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

16 March 2017 Version 1.08

Title word cross-reference

N [22].

-gram [22].

1.5 [220].

2.0 [216, 239, 274]. 2.0-inspired [274].

600 [76].

A&M [1]. ABC [44, 54]. Abuse [114]. Academic [266]. Access [10, 45, 55, 110, 131, 136, 185, 241, 243–245]. Accessible [186]. Acquiring [244]. Active [67, 69]. activity [269]. AD [76].

Adaptable [174]. Adaptation [202, 258]. Adaptations [128, 182]. Adaptive [119, 123, 143, 181, 183, 199–201, 203]. Addendum [134]. Adding [251]. Advanced [245]. Affecting [193]. Affordable [126, 246]. Age [89]. AgEcon [250]. Agency [172]. Agenda [73]. Aggregated [167]. AGROVOC [138]. Air [229]. Algorithm [175]. ALOCOM [209]. Ambient [171]. Analysing [211]. Analysis [17, 62, 224]. ANcient [76]. Ann [28]. annotated [261]. Anthropology [60]. Application [40, 48, 150, 184, 187]. Applications [16, 115, 138, 139, 158, 159, 268]. Appreciate [27]. Approach [159, 172, 205, 229, 269]. Approaches [15, 118, 233]. Architects [124]. Architectural [160].

Architectural-Level [160]. Architecture [7, 8]. Archival [263, 276]. Archive [60, 244, 248]. Archives [188, 210, 246]. Archiving [279]. AriadneTool [158]. Art [3, 106, 218]. art-resource [3]. Artifacts [165]. Asian [78]. Asking [97]. Aspects [169, 235]. Assessment [196, 248]. Associated [120, 134, 135]. Association [79]. Associations [131]. Atmospheric [64]. Audience [234]. audiovisual [284]. Audit [188]. Augmenting [35]. Author [42, 50, 255]. Author-generated [42, 50]. Authoring [154, 249]. Automated [180, 196]. Automatic [155, 198]. Automatically [213]. aware [254]. Awareness [39, 47]. AWE [266].

BAMBI [10]. Banner [37]. Banner-Blind [37]. Banners [37]. Barriers [267]. Based [20, 22, 39, 47, 67, 70, 120, 132, 134, 135, 137, 143, 175, 198, 203, 223, 253, 269, 278]. Baseline [42, 50]. Benefits [267]. Better [10, 285]. Between [33, 122, 237, 250]. Beyond [279]. Bibliography [105, 211]. Big [143, 152]. Bioinformatics [150]. Bite [103]. Blind [37]. Body [85]. Book [71, 115, 201]. Bookmarks [222]. Books [113]. Born [272]. Breast [103]. Breathe [87]. Bridging [185]. Bringing [154]. Broadcasts [186]. Brown [186]. Browser [124]. Browsing [10, 176]. Buddhism [78]. Buddhist [79]. Build [197, 233]. Building [125, 137, 190, 261, 275, 280]. Business [125].

Cafe [12]. California [168]. Can [98]. Candour [84]. Capabilities [259]. CAPM [128]. Capturing [178]. Case [166, 200, 242, 250, 275]. Case-Study [242]. Catalogue [191]. Categorization [175]. Category [172]. Cave [225]. CBETA [79]. Census [195]. Center [1]. Centered [57, 68, 123]. Centred [165]. Certification [191]. Challenges [120, 134, 135]. Chan [77]. Chan/Seon/Zen [77]. Change [172].

Changing [171, 265]. CHANT [76]. Characteristics [224]. Characterizing [213]. Checklist [188]. Chempound [274]. Child [28]. CHinese [75, 76, 79]. choosing [208]. CIDOC [242]. CIDOC-CRM [242]. cinematic [26]. Classification [57, 155]. CLIF [277]. Cloud [276, 281]. clouds [278]. Codes [57]. Coherence [25]. Collaboration [280]. Collaborative [17, 223, 252]. Collaborative-Filtering [223]. Collection [64, 67]. Collections [75, 164, 228, 242]. Combined [78]. Common [8, 142]. Common-Use [142]. Commons [284]. Communication [4]. Community [162, 275, 283]. Comparison [3, 122]. complex [268]. Compliant [192]. Compound [249]. comprehensive [76]. Computational [16, 67]. Computer [153, 171, 269]. computer-mediated [269]. Concept [137]. Concept-based [137]. Conceptual [128, 185]. Conditions [219]. CONDITOR1 [231]. Conference [51, 52]. Conflict [181]. confused [28]. Connection [58]. connections [34]. Connectivity [133]. Considerations [21]. Constraint [20]. Constraint-Based [20]. Construction [77, 269]. Content [69, 82, 130, 132, 155, 178, 182, 202, 209, 214, 262, 267, 277, 284]. Content-Based [132]. Context [56, 224]. Contribution [193]. Control [216]. Conversion [175, 200]. Cooperation [216]. Cooperative [18, 164]. Core [39, 42, 45, 47, 50–52, 55, 121, 242]. Corpora [173]. Corpus [261]. Cosmology [153]. Costs [126]. Counting [126]. Craft [82, 92]. Craft-Orientation [82]. Creation [164]. Creative [236]. Crisis [168]. Criteria [191, 196, 197]. Critic [86]. Critical [156]. Criticism [29, 81, 84, 85, 87, 89, 91, 92, 99, 101, 102, 104– 106]. Critics [100]. Critique [6]. CRM [242]. Crosses [170]. Cultural [239–241, 243]. Culture [83, 140]. Curation [194, 260, 264, 282]. Curatorial [275].

Current [39, 47]. Cycles [179].

DAR [271]. Dark [26]. Data [62, 119, 123, 153, 205, 240, 241, 248, 253, 274, 279, 282, 283, 285]. Database [3, 76, 77, 234]. Datuments [152]. DDC [261]. DDC-annotated [261]. Debates [245]. Declarative [38, 46]. delivering [284]. Deregulation [177]. Description [38, 46, 263]. Design [17, 111, 157, 158, 165, 166, 187, 244]. Designers [71]. Designing [16, 65, 174, 236]. Detecting [133]. Developing [120, 134, 135, 201, 210]. Development [14, 170]. Developments [64, 77]. Devices [182]. Dictionaries [78]. differences [250]. Different [153]. Digest [27]. Digital [1, 4, 19, 59–61, 65, 69, 75, 78, 79, 106, 107, 109, 118, 125, 126, 128–130, 161– 166, 169, 170, 176, 179, 185, 188, 191, 194, 196, 197, 201, 206, 210, 212, 214, 216, 234, 239, 242, 248, 252, 270, 272, 276, 281]. Digitization [127]. Disclosure [246]. Discovery [259]. Display [119, 123]. Dissemination [72]. Distance [115]. Distributed [62, 253]. DITA [231]. Diversity [264]. Divide [170]. Document [117, 155, 272]. Documents [144, 231]. Does [40, 48]. Domain [241]. Domain-Specific [241]. Domains [33]. DSpace [125, 272]. Dublin [39, 42, 45, 47, 50–52, 55, 242]. Dynamic [202].

E-Book [71, 115]. E-education [111, 116]. E-learning [183]. E-nhance [112]. E-notebook [120, 134, 135]. Early [218]. East [78]. Economic [93, 127, 130, 131]. Editing [249]. Editorial [5, 19, 29, 32, 51, 61, 75, 104, 105, 111, 130, 139, 142, 157, 163, 169]. Education [63, 67, 69, 111, 115, 116]. Educational [70, 165, 224]. Educators [167]. Efficiency [21]. Electric [177]. Electronic [12, 43, 53, 79, 96, 110, 113, 202]. embracing [266]. Emerging [128]. empirical [3]. Employee [172]. empower [256]. Enable [33, 241]. Enabling [72]. encouraging [283]. Encyclopaedias [202]. End [183, 254]. End-User [183]. Energy [168]. Engineering [67, 259]. Enhanced [221]. Enhancing [280]. Enron [168]. Enterprises [14]. Environment [203, 247, 264, 266]. Environmental [62]. Environments [68, 171, 270]. eResearch [254]. Escaping [90, 150]. Establishing [130]. Evaluating [128]. Evaluation [68, 111, 117, 180, 191, 198, 199]. Event [58]. Example [138, 268]. Exceed [267]. Excessive [84]. exchange [30]. Exhibits [178]. Existing [7]. Experience [36, 183, 215]. Experiences [167]. Experimental [68]. Experiments [218, 282]. expert [3]. Exploiting [239]. Exploratory [224]. Exploring [232]. Expressing [228]. Expression [175]. Extending [241].

Factors [130, 193]. Faculty [193]. FAME [201]. Fan [83]. Fast [159]. Fault [173]. Fault-tolerant [173]. Feature [156]. Federated [66]. Federations [252]. Fedora [284]. Feedback [213]. Fiction [83, 84]. Figures [77]. Files [272]. Filtering [223]. Finding [186]. Findings [195]. Físchlár [65]. FISHNet [283]. Flows [160]. fly [197]. Folksonomy [217, 218]. Formal [219]. Format [82]. forms [171]. Framework [72, 180, 184, 209, 217, 252, 258]. freshwater [283]. Fruit [279]. Fulltext [173]. Functionality [13, 15, 17, 78, 133]. Future [2, 142, 149, 163]. Futurology [108].

game [28]. Gap [185]. General [124, 234]. Generated [42, 50, 214]. Generic [200, 258]. Geographic [129]. Get [98]. Getting [95]. Goal [219]. Got [151]. Government [188]. gram [22]. grammars [198]. graphic [3]. Grassroots [94, 96]. Gray [57]. Grid [264]. GUI [124]. guide [208].

Handbooks [110]. Hands [95]. Hanging [279]. Having [34]. Here [98]. Heritage [239, 241, 243, 246]. heterogeneity [266]. Heterogeneous [66, 182]. Hillesund [74]. Historic [186]. Historical [176, 231]. Honing [92]. Horizontal [70]. human [171]. human-computer [171]. Hybrid [240, 278]. Hydra [280]. Hyperfiction [27]. Hypermedia [5–9, 16, 18, 68, 73, 100, 152, 158, 181, 199, 200]. Hypernews [25]. Hyperstructures [150]. Hypertext [13, 15, 17, 24, 26, 29, 36, 58, 70, 81, 91, 104, 105, 142, 149, 151, 154, 180, 225–230].

ICE [254]. ICE-Theorem [254]. ICT [177]. Identifiers [255]. Identifying [219]. If [197]. Images [10]. Impact [37, 127]. Impeding [193]. Implementation [120, 134, 135, 166, 184, 242]. Important [101]. Inappropriate [82]. Incorporating [62]. Indirect [154]. Industry [177]. Inevitable [87]. Influence [171, 215]. Informal [219]. Information [19, 21, 22, 68, 72, 89, 109, 121, 124, 127, 129, 146, 157, 160, 169, 172, 203, 207, 214, 220, 221, 231, 235, 236, 238, 243, 245, 268]. Infrastructure [146, 153, 254, 281]. ing [82]. Input [74]. Inputs [71]. Inquiry [67, 162]. Inquiry-based [67]. Insecurity [90]. inspired [274]. Institution [280]. Institutional [125, 192, 193, 195, 250, 265, 267, 271]. Integrating [144, 161, 284]. Integration [121]. Integrator [146]. Intelligence [171]. Intent [101]. interaction [171]. Interactional [59, 61]. Interactions [232]. Interactive [62]. Interactivity [59, 61]. Interchange [259]. Interdisciplinary [248]. Interface [3, 65, 66, 256]. Interfaces [180]. Intermediary [268]. Internet [149, 257]. Interoperability [5, 33, 199, 264]. Interpretations [228]. Introduction [13, 19, 29, 32, 51, 52, 59, 61, 75, 139, 142, 157, 169, 214, 230]. Involvement [166]. Issue [13, 19, 29, 32, 51, 52, 59, 61, 75, 139, 142, 157, 169, 214, 221, 230, 235, 243]. Issues [4, 19, 220]. Item [11].

Journal [214]. Journals [12]. Juliet [28]. Jus [12]. Just [87]. justified [122].

Kernel [43, 53]. Keyphrase [117]. keywords [211]. Kindura [278]. Knowledge [24, 30, 32, 57, 136, 139, 146, 244]. KOS [136].

Lab [1]. labelling [231]. Language

[38, 46, 160]. Large [22, 173]. Large-Scale [22]. Layers [175]. Layout [122]. layperson [256]. Lead [90]. Learning [83, 84, 111, 114, 137, 174, 183, 219, 221]. Lectures [112]. Left [122]. Left- [122]. Level [160]. LibQUAL [128]. Librarie [125]. Libraries [1, 59, 61, 69, 127, 130, 161, 163, 165, 185, 214, 239, 276]. Library [4, 45, 55, 65, 75, 128, 166, 176, 212, 216, 234, 281]. lifecycle [277]. Lightweight [156]. LIMBER [34]. lingual [224]. Link [147]. Linked [69]. Linking [56, 77]. Links [23, 26]. Literary [78, 249]. Literature [36, 94]. Long [248, 265, 276]. Long-Term [248, 276]. Longitudinal [234]. Lost [220]. Low [279].

Machine [40, 48, 153].

Machine-Understandable [40, 48]. Macro [118]. Magic [107]. Main [224]. Maintenance [156]. Make [189]. Making [256]. Management [18, 181, 207, 211, 235, 236, 253, 268, 285]. managers [262]. Managing [130, 155]. Manuscripts [10, 179]. Many [71, 74]. Map [70]. Map-Based [70]. Mapping [31, 141]. Maps [176, 231]. Markets [127]. Markup [179]. Martin [28]. MatDL [161]. Mathematical [67, 149]. Mathematics [149]. May [90]. Mechanisms [153].

MediaMosa [284]. mediated [269]. Mediates [168]. Medieval [179]. Membership [172]. Memory [20]. Menus [122]. Merging [132]. MERLOT [166]. Meta [72]. Metadata [33, 38, 40, 42, 43, 45, 46, 48, 50– 53, 55, 132, 167, 192, 208, 242, 261]. Metainformation [145]. MetaNet [33]. Metaphor [260]. METAXPath [41, 49]. Method [156]. Methodology [156]. Mexico [279]. Micro [260]. Micro-Services [260]. Migration [7]. Millenium [79]. Mining [233]. Mirror [93]. Misguided [88]. Model [44, 54, 153, 166, 181, 185, 187, 206, 219, 269]. model-based [269]. Modelling [160]. Models [157]. Modern [179, 271]. Modular [110]. Motivating [193]. Moving [78, 277]. Multi [181, 224, 280]. Multi-Institution [280]. Multi-lingual [224]. Multi-model [181]. Multilingual [45, 55, 173]. Multimedia [20]. Multimodal [182]. Murphy [186]. Museums [178, 218]. Music [228].

Names [11]. narration [26]. narrative [269]. Narratives [40, 48]. Nationwide [195]. nature [265]. Navigation [37, 70, 122]. Necessary [99, 244]. Need [103]. Needs [163, 244]. Negotiating [163]. nestor [191]. Netlib [67]. Network [60]. Networked [30, 32, 82]. Newspapers [176]. Next [143, 152]. nhance [112]. Nigerian [170]. Nightmare [80]. Nine [208]. No [103]. Non [57, 272]. Non-Born-Digital [272]. Non-reflected [57]. Northwest [210]. notebook [120, 134, 135]. novice [3]. NSDL [62, 64]. Numerical [175].

OAI [66, 192, 251, 261, 263]. OAI-Compliant [192]. Object [174, 242]. Objects [114, 249]. Ogun [170]. OHRA [7]. Oke [170]. One [74]. Ongoing [218]. Onion [175]. Ontologies [233]. Ontology [44, 54, 121]. oooxxxooo [28]. Open [5–9, 68, 131, 229, 247, 259, 276]. operating [250]. Optimal [57]. Options [127]. Order [233]. ORE [251, 263]. Organisation [136]. Organization [32, 42, 50, 57, 119, 123, 139, 238]. Organizational [171, 207]. Organizations [168]. Orientation [82]. Oriented [183, 258]. Other [227]. Out-of-turn [232]. Outputs [71, 74].

Page [122]. Paper [223]. Papers [51, 52, 223]. Paradigm [154]. Parliamentary [245]. Partnering [164]. past [2]. Path [7]. Patterns [133, 174]. Pedagogical [223]. Penn [275]. people [171]. Perception [37]. Performance [19, 22]. Permanence [43, 53]. Personal [146, 203, 207, 215, 228, 235, 236, 238]. Personalising [113]. Personalization [205]. Perspective [169, 258]. Phenomenology [109]. Philological [10]. physical [274]. Pieces [178]. Pipeline [260]. Plan [125]. platform [200]. Platforms [251]. Player [201]. Pliny [206]. PlumbingXJ [159]. Points [100]. Policy [253]. Policy-based [253]. Portal [167]. Positioning [127]. possibilities [35]. post [266]. Potential [115]. Power [164, 177]. practical [262]. Practice [161, 275]. Practitioners [236]. Pragmatics [23]. Preliminary [195]. present [2]. Preservation [126, 252, 265, 276]. Preserving [262, 284]. Pricing [129]. Principle [145]. principles [256]. Prison [150]. Problem [226]. Problems [4, 31, 77]. Process [258, 282]. Process-Oriented [258]. Processes [157]. Profiles [40, 48, 184]. Programmatic [136]. progressive [101]. Project [34, 170, 190]. Proposal [149]. Protocol [6]. Prototype [242]. Prototyping [159]. Provenance [282]. Provision [280]. proXimity [147]. Psychological [235]. Publications [257, 265]. Publishers [71]. Publishing [131]. Putting [240].

Quality [183]. questions [208].

RaaS [273]. RDD [38, 46]. RDF [30, 38, 46]. Re [129]. Re-visiting [129]. Reader [86, 213]. Reading [28, 36, 144]. Real [62, 199]. Real-time [62]. Recollection [237]. Recommender [219, 223, 259]. Records [285]. Recruiting [267]. REDDNET [276]. Redocumenting [269]. Reengineering [138]. Reference [7, 8]. Refinding [237]. reflected [57]. Related [134]. Relation [124]. Relational [77]. Relationship [237]. Relationships [35, 190, 252]. report [1]. Repositories [15, 66, 192–195, 255, 260, 264–266, 277]. Repository [125, 126, 188, 189, 191, 196, 197, 247, 248, 250, 251, 262, 267, 271, 273, 274, 278, 280]. representation [30]. Research [73, 118, 218, 268, 276, 285]. researchers [278]. Resistance [172]. resource [3]. Resources [42, 50, 162, 174, 224, 241]. response [74]. Restoring [252]. Restricted [270]. Retrieval [21, 22, 35, 132, 220, 221]. Reusable [114]. Reuse [209, 283]. Review [83, 207, 217]. Reviewing [102]. Rich [257]. Right [34, 122]. Right-justified [122]. Role [89]. Rural [170].

s [82]. Salience [213]. Scalability [20, 22, 271]. Scalable [21, 209]. Scale [22]. schema [208]. Schemas [184, 268]. Scholar [80]. Scholarly [4, 131, 255, 265]. Scholarship [206, 249]. Science [67, 84, 274, 283]. Scientific [127, 161, 173, 211, 248]. Search [88, 173, 180, 250]. searchers [3]. Searching [66, 118, 176]. Secondary [118]. Seeking [68]. Selected [51, 52, 105]. Self [248]. Self-Assessment [248]. Semantic [31, 73, 137, 140, 203, 239–241, 256]. semantically [254]. Semantics [242]. Sematic [33]. Seon [77]. Serial [11]. Servers [21]. Service [39, 47, 136, 172, 273]. Services [62, 128, 141, 189, 204, 212, 239, 260, 278, 279]. Sets [117]. Shadows [225]. Shared [205]. ShaRef [205]. Sharing [215, 283]. Sheer [282]. Sheet [20]. Show [257]. Site [122, 133]. sketch [101]. SMEs [156]. Social [169, 172, 217, 220–222, 235]. Software [14, 15, 67]. Solving [226]. Sonar [133]. Space [220, 238]. Spaces [137, 227]. Spatial [226–230]. Special [13, 19, 29, 32, 51, 52, 59, 61, 75, 139, 142, 157, 169, 214, 221, 230, 235, 243]. Specialized [233]. Specific [241]. Spoken [246]. Standards [5]. State [275]. Steps [194]. Storage [126, 276]. Strategic [127]. Strategies [238]. Structural [133]. Structure [58, 202]. Structures [207]. Student [238]. Study [1, 42, 50, 166, 170, 200, 234, 242, 250, 275]. Studying [217]. Style [20, 37]. subject [250]. Success [194, 197]. summary [198]. Support [94, 206, 251]. Supporting [14, 17, 162, 226]. Survey [12]. Sustaining [210]. Symbolic [175]. Syndication [259]. System [22, 68, 183, 187, 199, 244]. Systemless [229]. Systems [5, 7, 9, 18– 20, 32, 57, 136, 139, 143, 211, 219, 236, 253].

Tabular [119, 123]. Tagging [211, 215, 217, 218, 224]. Tags [224]. Talking [201]. Teaching [111]. Tech [276]. Technical [127]. Technique [175]. Techniques [66]. Technology [115, 172, 221]. Telescopes [63]. Television [186, 244]. Tell [257]. Term [33, 248, 265, 276]. Terminology [141]. Terms [40, 48, 78]. Texas [1, 276, 281]. Text [79, 149, 173, 175, 179, 198, 233]. Texts [76, 77, 79]. textual [3, 231]. Their [77, 171, 285]. Thematic [62]. Theorem [254]. There [98, 103]. Thesauri [138].

Thesaurus [31, 33, 35]. theses [254]. Thing [143, 152]. THREDDS [62]. time [62]. Together [280]. tolerant [173]. tool [231]. Toolkit [158]. Tools [62, 128, 137, 262]. Topic [231]. Trace [187]. traces [269]. transformations [225]. Tripitaka [79]. Trust [252]. Trusted [191, 194, 196]. Trustworthy [248]. Truth [88]. turn [232]. Twist [271]. Two [15]. Tyranny [90].

Ubiquitous [145]. UK [116]. ULIS [45, 55]. UML [30]. Understandable [40, 48]. Understanding [237]. Unified [150]. Universal [11]. Universe [153]. University [1, 276, 279]. Update [190]. upstream [277]. usability [3]. usable [256]. Usage [234]. Use [40, 48, 114, 142, 165, 192]. Usenet [72]. User [12, 57, 65, 68, 123, 163, 165, 166, 183, 211, 214, 219, 244]. User-Centered [57, 68, 123]. User-Centred [165]. User-Generated [214]. Users [37]. Using [30, 57, 167, 213].

Validation [196]. Value [223]. Valuing [129]. versus [102]. Viability [131]. Video [65]. Viewers [272]. Virtual [14, 63, 153, 178, 204]. Vision [9]. Visions [142]. visiting [129]. Visual [165, 226]. Visualising [222]. Visualization [64, 153]. Visualizing [249, 285]. Vocabulary [141]. vs [82].

Walking [147]. Warning [90]. Web [2, 37, 42, 50, 68, 73, 120, 122, 134, 135, 140, 143, 148, 154, 156, 159, 160, 187, 203, 204, 216, 220, 233, 239, 240, 256, 259, 274]. Web-Based [120, 134, 135, 143]. Websites [232]. Where [96]. Whistleblowing [270]. Wide [2]. WIED [160]. will [197]. Within [234, 252]. Work [134]. Workflow [24]. Working [266]. Workstation [10]. world [2, 199]. Wreader [27]. Writing [81, 104, 144, 148]. WYWWYWI [145].

XML [71, 74, 78, 184, 268].

You’ve [151].

Zen [77]. zetoc [39, 47].

References

Anonymous:1997:LRC

[1] Anonymous. Lab report: Center for the Study of Digital Libraries, Texas A&M University. Journal of Digital Information, 1(1):??, April 1997. CODEN ???? ISSN 1368-7506.

                                             URL http://jodi.ecs.soton.ac.uk/
                                             Articles/v01/i01/amdigli.

Berners-Lee:1996:WWW

[2] T. Berners-Lee. The World Wide Web — past, present and future. Journal of Digital Information, 1(1):??, July 1996. CODEN ???? ISSN 1368-7506.

                                             URL http://jodi.ecs.soton.ac.uk/
                                             Articles/v01/i01/BernersLee.

Dillon:1997:ECU

[3] A. Dillon and M. Song. An empirical comparison of the usability for novice and expert searchers of a textual and a graphic interface to an artresource database. Journal of Digital Information, 1(1):??, September 1997. CODEN ???? ISSN 1368-7506.

                                             URL http://jodi.ecs.soton.ac.uk/
                                             Articles/v01/i01/Dillon.

Harter:1997:SCD

[4] S. Harter. Scholarly communication and the digital library: Problems and issues. Journal of Digital Information, 1(1):??, April 1997. CODEN ???? ISSN 1368-

                                             7506. URL http://jodi.ecs.soton.
                                             ac.uk/Articles/v01/i01/Harter.

Wiil:1997:EOH

[5] U. Wiil. Editorial: Open hypermedia: Systems, interoperability and standards. Journal of Digital Information, 1(2):??, December 1997. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v01/i02/editorial.php.

Anderson:1997:COH

[6] K. Anderson, R. Taylor, and E. Whitehead, Jr. A critique of the Open Hypermedia Protocol. Journal of Digital Information, 1(2):??, December 1997. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v01/i02/Anderson.

Goose:1997:OTO

[7] S. Goose, A. Lewis, and H. Davis. OHRA: Towards an open hypermedia reference architecture and a migration path for existing systems. Journal of Digital Information, 1(2):??, December 1997. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v01/i02/Goose.

Gronbaek:1997:TCR

[8] K. Gronbaek and U. Wiil. Towards a common reference architecture for open hypermedia. Journal of Digital Information, 1(2):??, December 1997. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v01/i02/Gronbak.

Nurnberg:1997:VOH

[9] P. Nurnberg and J. Leggett. A vision for open hypermedia systems. Journal of Digital Information, 1(2):??, December 1997. CODEN ???? ISSN 1368-

                                        7506. URL http://jodi.ecs.soton.
                                        ac.uk/Articles/v01/i02/Nurnberg.

Calabretto:1998:PWB

[10] S. Calabretto and A. Bozzi. The philological workstation BAMBI (Better Access to Manuscripts and Browsing of Images). Journal of Digital Information, 1(3):??, September 1998. CODEN ???? ISSN 1368-7506.

                                        URL http://jodi.ecs.soton.ac.uk/
                                        Articles/v01/i03/Calabretto.

Cameron:1998:TUS

[11] R. Cameron. Towards universal serial item names. Journal of Digital Information, 1(3):??, August 1998. CODEN ???? ISSN 1368-7506.

                                        URL http://jodi.ecs.soton.ac.uk/
                                        Articles/v01/i03/Cameron.

Woodward:1998:CJE

[12] H. Woodward, F. Rowland, C. McKnight, C. Pritchett, and J. Meadows. Cafe Jus: an electronic journals user survey. Journal of Digital Information, 1(3):??, September 1998. CODEN ???? ISSN 1368-7506.

                                        URL http://jodi.ecs.soton.ac.uk/
                                        Articles/v01/i03/Woodward.

Ashman:1999:HFI

[13] H. Ashman, H. Oinas-Kukkonen, and M. Bieber. Hypertext functionality: introduction to the special issue. Journal of Digital Information, 1(4):??, February 1999. CODEN ???? ISSN 1368-

                                        7506. URL http://jodi.ecs.soton.
                                        ac.uk/Articles/v01/i04/editorial.

Noll:1999:SSD

[14] J. Noll and W. Scacchi. Supporting software development in virtual enterprises.

Journal of Digital Information, 1(4):??, January 1999. CODEN ???? ISSN 1368-

7506. URL http://jodi.ecs.soton.
ac.uk/Articles/v01/i04/Noll.

Oinas-Kukkonen:1999:TAS

[15] H. Oinas-Kukkonen and G. Rossi. On two approaches to software repositories and hypertext functionality. Journal of Digital Information, 1(4):??, January 1999. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v01/i04/Oinas-Kukkonen.

Rossi:1999:DCH

[16] G. Rossi, D. Schwabe, and A. Garrido. Designing computational hypermedia applications. Journal of Digital Information, 1(4):??, January 1999. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v01/i04/Rossi.

Selvin:1999:SCA

[17] A. Selvin. Supporting collaborative analysis and design with hypertext functionality. Journal of Digital Information, 1(4):??, January 1999. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v01/i04/Selvin.

Verbyla:1999:CHM

[18] J. Verbyla and C. Watters. Cooperative hypermedia management systems. Journal of Digital Information, 1(4):??, January 1999. CODEN ???? ISSN 1368-

7506. URL http://jodi.ecs.soton.
ac.uk/Articles/v01/i04/Verbyla.

Munson:2000:EPI

[19] E. Munson. Editorial: Performance issues in digital information systems: introduction to a special issue. Journal of Digital Information, 1(5):??, January 2000. CODEN ???? ISSN 1368-7506.

                                        URL http://jodi.ecs.soton.ac.uk/
                                        Articles/v01/i05/editorial.

Cumaranatunge:2000:MSC

[20] T. Cumaranatunge and E. Munson. Memory scalability in constraint-based multimedia style sheet systems. Journal of Digital Information, 1(5):??, January 2000. CODEN ???? ISSN 1368-7506.

                                        URL http://jodi.ecs.soton.ac.uk/
                                        Articles/v01/i05/Cumaranatunge.

Frieder:2000:ECS

[21] O. Frieder, D. Grossman, A. Chowdhury, and G. Frieder. Efficiency considerations for scalable information retrieval servers. Journal of Digital Information, 1(5):??, January 2000. CODEN ???? ISSN 1368-7506.

                                        URL http://jodi.ecs.soton.ac.uk/
                                        Articles/v01/i05/Frieder.

Miller:2000:PSL

[22] E. Miller, E. Shen, J. Liu, and C. Nicholas. Performance and scalability of a large-scale N-gram based information retrieval system. Journal of Digital Information, 1(5):??, January 2000. CODEN ???? ISSN 1368-7506.

                                        URL http://jodi.ecs.soton.ac.uk/
                                        Articles/v01/i05/Miller.

Tosca:2000:PL

[23] S. Tosca. A pragmatics of links. Journal of Digital Information, 1(6):??, June 2000. CODEN ???? ISSN 1368-7506.

                                        URL http://jodi.ecs.soton.ac.uk/
                                        Articles/v01/i06/Pajares.

Rada:2000:HKW

[24] R. Rada, A. Michailidis, C. Frosch, and M. Lei. Hypertext to knowledge to workflow. Journal of Digital Information, 1(6):??, May 2000. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v01/i06/Rada.

Engebretsen:2000:HC

[25] M. Engebretsen. Hypernews and coherence. Journal of Digital Information, 1(7):??, December 2000. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v01/i07/Engebretsen.

Miles:2000:HDC

[26] A. Miles. Hypertext in the dark: cinematic narration with links. Journal of Digital Information, 1(7):??, December 2000. CODEN ???? ISSN 1368-

7506. URL http://jodi.ecs.soton.
ac.uk/Articles/v01/i07/Miles.

Rau:2000:WDH

[27] A. Rau. Wreader’s digest — how to appreciate hyperfiction. Journal of Digital Information, 1(7):??, December 2000. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v01/i07/Rau.

Walker:2000:CGC

[28] J. Walker. Child’s game confused: reading Juliet Ann Martin’s oooxxxooo. Journal of Digital Information, 1(7): ??, December 2000. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v01/i07/
Walker.

Tosca:2001:EHC

[29] S. Tosca. Editorial: Hypertext criticism: introduction to a special issue. Journal of Digital Information, 1(7):??, January 2001. CODEN ???? ISSN 1368-

7506. URL http://jodi.ecs.soton.
ac.uk/Articles/v01/i07/editoria.

Cranefield:2001:NKR

[30] S. Cranefield. Networked knowledge representation and exchange using UML and RDF. Journal of Digital Information, 1(8):??, February 2001. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v01/i08/Cranefield.

Doerr:2001:SPT

[31] M. Doerr. Semantic problems of thesaurus mapping. Journal of Digital Information, 1(8):??, March 2001. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v01/i08/Doerr.

Hill:2001:ENK

[32] L. Hill and T. Koch. Editorial: Networked knowledge organization systems: introduction to a special issue. Journal of Digital Information, 1(8):??, April 2001. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v01/i08/editorial.

Hunter:2001:MMT

[33] J. Hunter. MetaNet — A metadata term thesaurus to enable sematic interoperability between metadata domains. Journal of Digital Information, 1(8):??, February 2001. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v01/i08/Hunter.

Miller:2001:HRC

[34] K. Miller and B. Matthews. Having the right connections: the LIM-BER project. Journal of Digital Information, 1(8):??, February 2001. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v01/i08/Miller.

Tudhope:2001:ATR

[35] D. Tudhope, H. Alani, and C. Jones. Augmenting thesaurus relationships: possibilities for retrieval. Journal of Digital Information, 1(8):??, February 2001. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v01/i08/Tudhope.

Miall:2001:RHE

[36] David S. Miall and Teresa Dobson. Reading hypertext and the experience of literature. Journal of Digital Information, 2(1):??, August 2001. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v02/i01/Miall;   http:/
/journals.tdl.org/jodi/article/
view/35;   http://journals.tdl.
org/jodi/article/view/35/37.

Pagendarm:2001:WUB

[37] Magnus Pagendarm and Heike Schaumburg. Why are users banner-blind? The impact of navigation style on the perception of Web banners. Journal of Digital Information, 2(1): ??, August 2001. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v02/i01/
Pagendarm;   http://journals.tdl.
org/jodi/article/view/36;   http:
//journals.tdl.org/jodi/article/
view/36/38.

Anutariya:2001:RDD

[38] C. Anutariya, V. Wuwongse, K. Akama, and E. Nantajeewarawat. RDF declarative description (RDD): A language for metadata. Journal of Digital Information, 2(2):??, November 2001. CODEN ???? ISSN 1368-7506.

                                        URL http://jodi.ecs.soton.ac.uk/
                                        Articles/v02/i02/Anutariya.

Apps:2001:ZDC

[39] A. Apps and R. MacIntyre. zetoc: a Dublin Core based current awareness service. Journal of Digital Information, 2(2):??, November 2001. CODEN ???? ISSN 1368-7506.

                                        URL http://jodi.ecs.soton.ac.uk/
                                        Articles/v02/i02/Apps.

Baker:2001:WTD

[40] T. Baker, M. Dekkers, R. Heery, M. Patel, and G. Salokhe. What terms does your metadata use? application profiles as machine-understandable narratives. Journal of Digital Information, 2(2):??, November 2001. CODEN ???? ISSN 1368-7506.

                                        URL http://jodi.ecs.soton.ac.uk/
                                        Articles/v02/i02/Baker.

Dyreson:2001:M

[41] C. Dyreson, M. Bohlen, and C. Jensen. MetaXPath. Journal of Digital Information, 2(2):??, November 2001. CODEN ???? ISSN 1368-7506.

                                        URL http://jodi.ecs.soton.ac.uk/
                                        Articles/v02/i02/Dyreson.

Greenberg:2001:AGD

[42] J. Greenberg, M. Pattuelli, B. Parsia, and W. Robertson. Authorgenerated Dublin Core Metadata for Web resources: A baseline study in an organization. Journal of Digital Information, 2(2):??, November 2001. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v02/i02/Greenberg.

Kunze:2001:MKE

[43] J. Kunze. A metadata kernel for electronic permanence. Journal of Digital Information, 2(2):??, November 2001. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v02/i02/Kunze.

Lagoze:2001:AOM

[44] C. Lagoze and J. Hunter. The ABC ontology and model. Journal of Digital Information, 2(2):??, November 2001. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v02/i02/Lagoze.

Wen:2001:MAD

[45] D. Wen, T. Sakaguchi, S. Sugimoto, and K. Tabata. Multilingual access to Dublin Core metadata of ULIS Library. Journal of Digital Information, 2(2):??, November 2001. CODEN ???? ISSN 1368-

7506. URL http://jodi.ecs.soton.
ac.uk/Articles/v02/i02/Wen.

Anutariya:2002:RDD

[46] Chutiporn Anutariya, Vilas Wuwongse, Kiyoshi Akama, and Ekawit Nantajeewarawat. RDF declarative description (RDD): A language for metadata. Journal of Digital Information, 2(2): ??, 2002. CODEN ???? ISSN 1368-

7506. URL http://journals.tdl.
org/jodi/article/view/38;   http:
//journals.tdl.org/jodi/article/
view/38/41.

Apps:2002:ZDC

[47] Ann Apps and Ross MacIntyre. zetoc: a Dublin Core based current awareness service. Journal of Digital Information, 2(2):??, 2002. CODEN ???? ISSN 1368-

7506. URL http://journals.tdl.
org/jodi/article/view/39;   http:
//journals.tdl.org/jodi/article/
view/39/42.

Baker:2002:WTD

[48] Thomas Baker, Makx Dekkers, Rachel Heery, Manjula Patel, and Gauri Salokhe. What terms does your metadata use? application profiles as machineunderstandable narratives. Journal of Digital Information, 2(2):??, 2002. CODEN ???? ISSN 1368-

7506. URL http://journals.tdl.
org/jodi/article/view/40;   http:
//journals.tdl.org/jodi/article/
view/40/43.

Dyreson:2002:M

[49] Curtis E. Dyreson, Michael H. Bohlen, and Christian S. Jensen. METAXPath. Journal of Digital Information, 2(2): ??, 2002. CODEN ???? ISSN 1368-

7506. URL http://journals.tdl.
org/jodi/article/view/41;   http:
//journals.tdl.org/jodi/article/
view/41/44.

Greenberg:2002:AGD

[50] Jane Greenberg, Maria Cristina Pattuelli, Bijan Parsia, and W. Davenport Robertson. Author-generated Dublin Core metadata for Web resources: A baseline study in an organization. Journal of Digital Information, 2(2):??, 2002. CODEN ???? ISSN 1368-

7506. URL http://journals.tdl.
org/jodi/article/view/42;   http:
//journals.tdl.org/jodi/article/
view/42/45.

Koch:2002:EIS

[51] T. Koch and S. Weibel. Editorial: Introduction to a special issue on metadata: Selected papers from the Dublin Core 2001 Conference. Journal of Digital Information, 2(2):??, January 2002. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v02/i02/editorial.

Koch:2002:ISI

[52] Traugott Koch and Stuart Weibel. Introduction to a special issue on metadata: Selected papers from the Dublin Core 2001 Conference. Journal of Digital Information, 2(2):??, 2002. CODEN ???? ISSN 1368-

7506. URL http://journals.tdl.
org/jodi/article/view/37/39.

Kunze:2002:MKE

[53] John A. Kunze. A metadata kernel for electronic permanence. Journal of Digital Information, 2(2):??, 2002. CODEN ???? ISSN 1368-

7506. URL http://journals.tdl.
org/jodi/article/view/43;   http:
//journals.tdl.org/jodi/article/
view/43/46.

Lagoze:2002:AOM

[54] Carl Lagoze and Jane Hunter. The ABC ontology and model. Journal of Digital Information, 2(2):??, 2002. CODEN ???? ISSN 1368-

7506. URL http://journals.tdl.
org/jodi/article/view/44;   http:
//journals.tdl.org/jodi/article/
view/44/47.

Wen:2002:MAD

[55] Danyang Wen, Tetsuo Sakaguchi, Shigeo Sugimoto, and Koichi Tabata.

Multilingual access to Dublin Core metadata of ULIS library. Journal of Digital Information, 2(2):??, 2002. CODEN ???? ISSN 1368-

7506. URL http://journals.tdl.
org/jodi/article/view/45;   http:
//journals.tdl.org/jodi/article/
view/45/48.

El-Beltagy:2002:LC

[56] Samhaa R. El-Beltagy, Wendy Hall, David De Roure, and Leslie Carr. Linking in context. Journal of Digital Information, 2(3):??, March

2002.  CODEN ????  ISSN

1368-7506. URL http://jodi.ecs.
soton.ac.uk/Articles/v02/i03/El-
Beltagy;   http://journals.tdl.
org/jodi/article/view/46;   http:
//journals.tdl.org/jodi/article/
view/46/49.

Losee:2002:OUC

[57] Robert M. Losee. Optimal user-centered knowledge organization and classification systems: Using non-reflected gray codes. Journal of Digital Information, 2(3):??, February 2002. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v02/i03/Losee;   http:/
/journals.tdl.org/jodi/article/
view/47;   http://journals.tdl.
org/jodi/article/view/47/50.

Miles:2002:HSE

[58] Adrian Miles. Hypertext structure as the event of connection. Journal of Digital Information, 2(3):??, March 2002. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v02/i03/Miles;   http:/
/journals.tdl.org/jodi/article/

view/48;   http://journals.tdl.
org/jodi/article/view/48/51.

Admin:2002:IDL

[59] Anita Admin and Maliaca Oxnam. Interactional digital libraries: introduction to a special issue on interactivity in digital libraries. Journal of Digital Information, 2(4):??, 2002. CODEN ???? ISSN 1368-

7506. URL http://journals.tdl.
org/jodi/article/view/49/52.

Clark:2002:DAN

[60] Jeffrey T. Clark, Brian M. Slator, James E. Landrum III, Richard Frovarp, Aaron Bergstrom, Sanjay Ramaswamy, and William Jockheck. Digital archive network for anthropology. Journal of Digital Information, 2(4):??, May 2002. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v02/i04/Clark;   http:/
/journals.tdl.org/jodi/article/
view/50;   http://journals.tdl.
org/jodi/article/view/50/53.

Coleman:2002:EID

[61] A. Coleman and M. Oxnam. Editorial: Interactional digital libraries: Introduction to the special issue on interactivity in digital libraries. Journal of Digital Information, 2(4):??, May 2002. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v02/i04/editoria.

Domenico:2002:TRT

[62] Ben Domenico, John Caron, Ethan Davis, Robb Kambic, and Stefano Nativi. Thematic Real-time Environmental Distributed Data Services (THREDDS): Incorporating interactive analysis tools into NSDL. Journal of Digital Information, 2(4):??, May 2002. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v02/i04/
Domenico;   http://journals.tdl.
org/jodi/article/view/51;   http:
//journals.tdl.org/jodi/article/
view/51/54.

Hoban:2002:VTE

[63] Susan Hoban, Marie desJardins, Nora Farrell, Priyang Rathod, Joel Sachs, Suryakant Sansare, Yelena Yesha, John Keating, Bart Busschots, Johanna Means, Gilbert Clark, Louuis Mayo, and Willard Smith. Virtual telescopes in education. Journal of Digital Information, 2(4):??, May 2002. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v02/i04/Hoban;   http:/
/journals.tdl.org/jodi/article/
view/52;   http://journals.tdl.
org/jodi/article/view/52/55.

Klaus:2002:AVC

[64] Christopher Klaus, Keith Andrew, and Gerald Mace. Atmospheric visualization collection: Developments in the NSDL. Journal of Digital Information, 2(4):??, May 2002. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v02/i04/Klaus;   http:/
/journals.tdl.org/jodi/article/
view/53;   http://journals.tdl.
org/jodi/article/view/53/56.

Lee:2002:DUI

[65] Hyowon Lee and Alan F. Smeaton. Designing the user interface for the Físchlár Digital Video Library. Journal of Digital Information, 2(4):??, May 2002. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v02/i04/
Lee;   http://journals.tdl.org/
jodi/article/view/54;   http://
journals.tdl.org/jodi/article/view/
54/57.

Liu:2002:FSI

[66] Xiaoming Liu, Kurt Maly, Mohammad Zubair, Qiaoling Hong, Michael L. Nelson, Frances Knudson, and Irma Holtkamp. Federated searching interface techniques for heterogeneous OAI repositories. Journal of Digital Information, 2(4):??, May 2002. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v02/i04/
Liu;   http://journals.tdl.org/
jodi/article/view/55;   http://
journals.tdl.org/jodi/article/view/
55/58.

Moore:2002:ANA

[67] Shirley Moore, A. J. Baker, Jack Dongarra, Christian Halloy, and Chung Ng. Active Netlib: An active mathematical software collection for inquirybased computational science and engineering education. Journal of Digital Information, 2(4):??, May 2002. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v02/i04/Moore;   http:/
/journals.tdl.org/jodi/article/
view/56;   http://journals.tdl.
org/jodi/article/view/56/59.

Salampasis:2002:EUC

[68] Michail Salampasis and Konstantinos I. Diamantaras. Experimental user-centered evaluation of an open hypermedia system and Web information seeking environments. Journal of Digital Information, 2(4):??, May 2002. CODEN ???? ISSN 1368-7506. URL http://jodi.

                                        ecs.soton.ac.uk/Articles/v02/i04/
                                        Salampasis;   http://journals.tdl.
                                        org/jodi/article/view/57;   http:
                                        //journals.tdl.org/jodi/article/
                                        view/57/60.

Yaron:2002:LAC

[69] David Yaron, D. Jeff Milton, and Rea Freeland. Linked active content for digital libraries for education. Journal of Digital Information, 2(4):??, May 2002. CODEN ???? ISSN 1368-7506.

                                        URL http://jodi.ecs.soton.ac.uk/
                                        Articles/v02/i04/Yaron;   http:/
                                        /journals.tdl.org/jodi/article/
                                        view/58;   http://journals.tdl.
                                        org/jodi/article/view/58/63.

Brusilovsky:2002:MBH

[70] Peter Brusilovsky and Riccardo Rizzo. Map-based horizontal navigation in educational hypertext. Journal of Digital Information, 3(1):??, July 2002. CODEN ???? ISSN 1368-7506. URL http://jodi.

                                        ecs.soton.ac.uk/Articles/v03/i01/
                                        Brusilovsky;   http://journals.
                                        tdl.org/jodi/article/view/75;
                                        http://journals.tdl.org/jodi/article/
                                        view/75/74.

Hillesund:2002:MOM

[71] Terje Hillesund. Many outputs — many inputs: XML for publishers and E-book designers. Journal of Digital Information, 3(1): ??, August 2002. CODEN ???? ISSN 1368-7506. URL http://jodi.

                                        ecs.soton.ac.uk/Articles/v03/i01/

Hillesund;   http://journals.tdl.
org/jodi/article/view/76;   http:
//journals.tdl.org/jodi/article/
view/76/75.

Lueg:2002:EDM

[72] Christopher Lueg. Enabling dissemination of meta information in the Usenet framework. Journal of Digital Information, 3(1):??, 2002. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v03/i01/Lueg;   http:/
/journals.tdl.org/jodi/article/
view/77;   http://journals.tdl.
org/jodi/article/view/77/76.

vanOssenbruggen:2002:HSW

[73] Jacco van Ossenbruggen, Lynda Hardman, and Lloyd Rutledge. Hypermedia and the Semantic Web: A research agenda. Journal of Digital Information, 3(1):??, May 2002. CODEN ???? ISSN 1368-7506. URL http:

//jodi.ecs.soton.ac.uk/Articles/
v03/i01/VanOssenbruggen;   http:
//journals.tdl.org/jodi/article/
view/78;   http://journals.tdl.
org/jodi/article/view/78/77.

Walsh:2002:XOI

[74] Norman Walsh. XML: One input — many outputs: a response to Hillesund. Journal of Digital Information, 3(1):??, September 2002. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v03/i01/Walsh;   http:/
/journals.tdl.org/jodi/article/
view/79;   http://journals.tdl.
org/jodi/article/view/79/78.

Bruya:2002:ECC

[75] Brian Bruya. Editorial: Chinese collections in the digital library: introduction to a special issue. Journal of Digital Information, 3(2): ??, October 2002. CODEN ???? ISSN 1368-7506. URL http://jodi.

                                        ecs.soton.ac.uk/Articles/v03/i02/
                                        editorial;   http://journals.tdl.
                                        org/jodi/article/view/80/79.

Ho:2002:CCA

[76] Che Wah Ho. CHANT (CHinese ANcient Texts): a comprehensive database of all ancient Chinese texts up to 600 AD. Journal of Digital Information, 3 (2):??, August 2002. CODEN ???? ISSN

                                        1368-7506. URL http://jodi.ecs.
                                        soton.ac.uk/Articles/v03/i02/Ho;
                                         http://journals.tdl.org/jodi/article/
                                        view/81;   http://journals.tdl.
                                        org/jodi/article/view/81/80.

Mohr:2002:LCS

[77] Michel Mohr. Linking Chan/Seon/ Zen figures and their texts: Problems and developments in the construction of a relational database. Journal of Digital Information, 3(2):??, August 2002. CODEN ???? ISSN 1368-7506.

                                        URL http://jodi.ecs.soton.ac.uk/
                                        Articles/v03/i02/Mohr;   http:/
                                        /journals.tdl.org/jodi/article/
                                        view/82;   http://journals.tdl.
                                        org/jodi/article/view/82/81.

Muller:2002:MXF

[78] Charles Muller and Michael Beddow. Moving into XML functionality: The combined digital dictionaries of Buddhism and East Asian literary terms. Journal of Digital Information, 3(2):??, August 2002. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v03/i02/Muller;   http:
//journals.tdl.org/jodi/article/
view/83;   http://journals.tdl.
org/jodi/article/view/83/82.

Wittern:2002:CBT

[79] Christian Wittern. Chinese Buddhist texts for the new Millenium — the Chinese Buddhist Electronic Text Association (CBETA) and its Digital Tripitaka. Journal of Digital Information, 3(2):??, September 2002. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i02/
Wittern;   http://journals.tdl.
org/jodi/article/view/84;   http:
//journals.tdl.org/jodi/article/
view/84/83.

Higgason:2002:SN

[80] R. Higgason. A scholar’s nightmare. Journal of Digital Information, 3(3):??, January 2002. CODEN ???? ISSN 1368-

7506. URL http://jodi.ecs.soton.
ac.uk/Articles/v03/i03/Higgason/
nightmare.htm.

Anonymous:2003:HCW

[81] Anonymous. Hypertext criticism: Writing about hypertext. Journal of Digital Information, 3(3):??, 2003. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/117;   http://journals.tdl.
org/jodi/article/view/117/116.

Breeze:2003:IFC

[82] M. Breeze. Inappropriate format][ing][: Craft-orientation vs. networked content[s]. Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

                                        ecs.soton.ac.uk/Articles/v03/i03/
                                        Breeze/craft_or_network.htm.

Chatelain:2003:LRC

[83] J. Chatelain. Learning from the review culture of fan fiction. Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-

                                        7506. URL http://jodi.ecs.soton.
                                        ac.uk/Articles/v03/i03/Chatelain/
                                        fanfic.htm.

Chatelain:2003:LSF

[84] J. Chatelain. Learning from science fiction criticism: Excessive candour. Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-

                                        7506. URL http://jodi.ecs.soton.
                                        ac.uk/Articles/v03/i03/Chatelain/
                                        scifi.htm.

Higgason:2003:BC

[85] R. Higgason. A body of criticism. Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-

                                        7506. URL http://jodi.ecs.soton.
                                        ac.uk/Articles/v03/i03/Higgason/
                                        critbody.htm.

Higgason:2003:CR

[86] R. Higgason. Critic as reader. Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-

                                        7506. URL http://jodi.ecs.soton.
                                        ac.uk/Articles/v03/i03/Higgason/
                                        critread.htm.

Higgason:2003:MCI

[87] R. Higgason. Is more criticism inevitable? just breathe. Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-

7506. URL http://jodi.ecs.soton.
ac.uk/Articles/v03/i03/Higgason/
inevitable.htm.

Higgason:2003:MST

[88] R. Higgason. Misguided search for truth. Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i03/
Higgason/truth.htm.

Higgason:2003:RCI

[89] R. Higgason. A role of criticism in the information age. Journal of Digital Information, 3(3):??, January 2003. CO-DEN ???? ISSN 1368-7506. URL http:

//jodi.ecs.soton.ac.uk/Articles/
v03/i03/Higgason/infoage.htm.

Higgason:2003:WET

[90] R. Higgason. Warning: Escaping tyranny may lead to insecurity. Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v03/i03/Higgason/unread.
htm.

Higgason:2003:WHC

[91] R. Higgason. Why hypertext criticism? Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i03/
Higgason/whycrit.htm.

Larsen:2003:CHC

[92] D. Larsen. Criticism: Honing the craft. Journal of Digital Information, 3(3):??, January 2003. CODEN ????

ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i03/
Larsen/criticism.htm.

Larsen:2003:EM

[93] D. Larsen. The economic mirror. Journal of Digital Information, 3(3): ??, January 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i03/
Larsen/economics.htm.

Larsen:2003:GSL

[94] D. Larsen. Grassroots support literature. Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i03/
Larsen/linearroots.htm.

Larsen:2003:GYH

[95] D. Larsen. Getting your hands on it. Journal of Digital Information, 3 (3):??, January 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i03/
Larsen/access.htm.

Larsen:2003:WEG

[96] D. Larsen. Where are the electronic grassroots? Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-7506. URL http:

//jodi.ecs.soton.ac.uk/Articles/
v03/i03/Larsen/electronicroots.
htm.

Larsen:2003:WWA

[97] D. Larsen. What are we asking for? Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i03/
Larsen/readers.htm.

Larsen:2003:YCG

[98] D. Larsen. You can get there from here. Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i03/
Larsen/smallsteps.htm.

Marsh:2003:NCN

[99] B. Marsh. New criticism necessary? Journal of Digital Information, 3(3): ??, January 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i03/
Marsh/newcriticism.htm.

Marsh:2003:PHC

[100] B. Marsh. Points for hypermedia critics. Journal of Digital Information, 3 (3):??, January 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i03/
Marsh/pointsforcritics.htm.

Miles:2003:IIS

[101] A. Miles. Intent is important (a sketch for a progressive criticism). Journal of Digital Information, 3(3): ??, January 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i03/
Miles/intent.htm.

Miles:2003:RVC

[102] A. Miles. Reviewing versus criticism. Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i03/
Miles/reviewing&criticism.htm.

Miles:2003:TNN

[103] A. Miles. There’s no need to bite the breast. Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i03/
Miles/breast.htm.

Tosca:2003:EHC

[104] Susana Tosca and Jill Walker. Editorial: Hypertext criticism: Writing about hypertext. Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/116/115;   http://jodi.ecs.
soton.ac.uk/Articles/v03/i03/editorial.
htm.

Tosca:2003:ESB

[105] Susana Tosca and Jill Walker. Editorial: Selected bibliography of hypertext criticism. Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/115/114;   http://jodi.ecs.
soton.ac.uk/Articles/v03/i03/bibliography.
htm.

Weight:2003:DAC

[106] J. Weight. Digital art criticism. Journal of Digital Information, 3(3): ??, January 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i03/
Weight/digitalartcriticism.htm.

Weight:2003:DM

[107] J. Weight. Digital magic. Journal of Digital Information, 3(3): ??, January 2003. CODEN ????

ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i03/
Weight/digitalmagic.htm.

Weight:2003:F

[108] J. Weight. Futurology. Journal of Digital Information, 3(3):??, January 2003. CO-DEN ???? ISSN 1368-7506. URL http:

//jodi.ecs.soton.ac.uk/Articles/
v03/i03/Weight/futurology.htm.

Weight:2003:PDI

[109] J. Weight. Phenomenology and digital information. Journal of Digital Information, 3(3):??, January 2003. CODEN ???? ISSN 1368-7506. URL http:

//jodi.ecs.soton.ac.uk/Articles/
v03/i03/Weight/phenomenology.htm.

Caracciolo:2003:TMA

[110] Caterina Caracciolo. Towards modular access to electronic handbooks. Journal of Digital Information, 3(4): ??, February 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i04/
Caracciolo;   http://journals.tdl.
org/jodi/article/view/86;   http:
//journals.tdl.org/jodi/article/
view/86/84.

Landoni:2003:EED

[111] Monica Landoni and Paloma Diaz.

Editorial: E-education: Design and evaluation for teaching and learning. Journal of Digital Information, 3(4):??, March 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i04/
editorial;   http://journals.tdl.
org/jodi/article/view/118/85.

Naber:2003:NL

[112] Larissa Naber and Monika Köhle. Enhance lectures. Journal of Digital Information, 3(4):??, February 2003. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v03/i04/Naber;   http:/
/journals.tdl.org/jodi/article/
view/87;   http://journals.tdl.
org/jodi/article/view/87/86.

Ohene-Djan:2003:PEB

[113] James Ohene-Djan and Alvaro A. A. Fernandes. Personalising electronic books. Journal of Digital Information, 3(4):??, February 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i04/
Ohene-Djan;   http://journals.tdl.
org/jodi/article/view/88;   http:
//journals.tdl.org/jodi/article/
view/88/87.

Polsani:2003:UAR

[114] Pithamber R. Polsani. Use and abuse of reusable learning objects. Journal of Digital Information, 3(4):??, February 2003. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v03/i04/Polsani;   http:
//journals.tdl.org/jodi/article/
view/89;   http://journals.tdl.
org/jodi/article/view/89/88.

Shiratuddin:2003:BTP

[115] Norshuhada Shiratuddin, Monica Landoni, Forbes Gibb, and Shahizan Hassan. E-book technology and its potential applications in distance education. Journal of Digital Information, 3(4):??, February 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v03/i04/

Shiratuddin;   http://journals.
tdl.org/jodi/article/view/90;
http://journals.tdl.org/jodi/article/
view/90/89.

Wilson:2003:EU

[116] Ruth Wilson. E-education in the UK. Journal of Digital Information, 3(4):??, February 2003. CO-DEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v03/i04/Wilson;   http:
//journals.tdl.org/jodi/article/
view/91;   http://journals.tdl.
org/jodi/article/view/91/90.

Jones:2003:EDK

[117] Steve Jones and Gordon W. Paynter. An evaluation of document keyphrase sets. Journal of Digital Information, 4(1):??, February 2003. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v04/i01/Jones;   http:/
/journals.tdl.org/jodi/article/
view/93;   http://journals.tdl.
org/jodi/article/view/93/92.

Levitt:2003:MAD

[118] Jonathan Mendel Levitt. Macro approaches to digital searching and secondary research. Journal of Digital Information, 4(1):??, February 2003. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v04/i01/Levitt;   http:
//journals.tdl.org/jodi/article/
view/95;   http://journals.tdl.
org/jodi/article/view/95/94.

Losee:2003:AOT

[119] Robert M. Losee. Adaptive organization of tabular data for display.

Journal of Digital Information, 4(1): ??, 2003. CODEN ???? ISSN 1368-

7506. URL http://journals.tdl.
org/jodi/article/view/96;   http:
//journals.tdl.org/jodi/article/
view/96/95.

Reimer:2003:ICA

[120] Y. Jacobs Reimer and S. Douglas. Implementation challenges associated with developing a Web-based E-notebook. Journal of Digital Information, 4(3): ??, September 2003. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v04/i03/
JacobsReimer.

Doerr:2004:TCO

[121] Martin Doerr, Jane Hunter, and Carl Lagoze. Towards a core ontology for information integration. Journal of Digital Information, 4(1):??, April 2004. CODEN ???? ISSN 1368-7506.

URL http://jodi.ecs.soton.ac.uk/
Articles/v04/i01/Doerr;   http:/
/journals.tdl.org/jodi/article/
view/92;   http://journals.tdl.
org/jodi/article/view/92/91.

Kalbach:2003:WPL

[122] James Kalbach and Tim Bosenick. Web page layout: A comparison between left- and right-justified site navigation menus. Journal of Digital Information, 4(1):??, April 2004. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v04/i01/
Kalbach;   http://journals.tdl.
org/jodi/article/view/94;   http:
//journals.tdl.org/jodi/article/
view/94/93.

Losee:2003:AUC

[123] R. Losee. Adaptive user-centered organization of tabular data for display. Journal of Digital Information, 4(1):??, April 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.ecs.soton.
ac.uk/Articles/v04/i01/Losee.

Marchionini:2003:TGR

[124] Gary Marchionini and Ben Brunk.

Towards a general relation browser: A GUI for information architects. Journal of Digital Information, 4 (1):??, April 2004. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v04/i01/
Marchionini;   http://journals.
tdl.org/jodi/article/view/97;
http://journals.tdl.org/jodi/article/
view/97/96.

Barton:2004:BBP

[125] Mary R. Barton and Julie Harford Walker. Building a business plan for DSpace, MIT libraries’ digital institutional repository. Journal of Digital Information, 4(2):??, May 2004. CODEN ???? ISSN 1368-

7506. URL http://journals.tdl.
org/jodi/article/view/99;   http:
//journals.tdl.org/jodi/article/
view/99/98;   http://jodi.ecs.soton.
ac.uk/Articles/v04/i02/Barton.

Chapman:2004:CCD

[126] Stephen Chapman. Counting the costs of digital preservation: Is repository storage affordable? Journal of Digital Information, 4(2):??, May 2004. CO-DEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/100;   http://journals.tdl.
org/jodi/article/view/100/99;

http://jodi.ecs.soton.ac.uk/Articles/
v04/i02/Chapman.

Geyer-Schulz:2004:SPO

[127] Andreas Geyer-Schulz, Andreas Neumann, Annika Heitmann, and Karsten Stroborn. Strategic positioning options for scientific libraries in markets of scientific and technical information — the economic impact of digitization. Journal of Digital Information, 4(2):??, May 2004. CODEN ???? ISSN 1368-7506. URL http://jodi.

ecs.soton.ac.uk/Articles/v04/i02/
Geyer-Schulz;   http://journals.
tdl.org/jodi/article/view/101;
http://journals.tdl.org/jodi/article/
view/101/100.

Heath:2004:ETE

[128] Fred Heath, Martha Kyrillidou, Duane Webster, Sayeed Choudhury, Ben Hobbs, Mark Lorie, and Nicholas Flores. Emerging tools for evaluating digital library services: Conceptual adaptations of LibQUAL+ and CAPM. Journal of Digital Information, 4(2):??, June 2004. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/102;   http://journals.tdl.
org/jodi/article/view/102/101;
http://jodi.ecs.soton.ac.uk/Articles/
v04/i02/Heath.

Longhorn:2004:RVV

[129] Roger Longhorn and Michael Blakemore. Re-visiting the valuing and pricing of digital geographic information. Journal of Digital Information, 4(2):??, 2004. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/

view/103;   http://journals.tdl.
org/jodi/article/view/103/102.

Tanner:2004:EEF

[130] Simon Tanner. Editorial: Economic factors of managing digital content and establishing digital libraries. Journal of Digital Information, 4(2):??, June 2004. CODEN ???? ISSN

1368-7506. URL http://journals.
tdl.org/jodi/article/view/98/97;
http://jodi.ecs.soton.ac.uk/Articles/
v04/i02/editorial.

Willinsky:2004:SAE

[131] John Willinsky. Scholarly associations and the economic viability of open access publishing. Journal of Digital Information, 4(2):??, April 2004. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/104;   http://journals.tdl.
org/jodi/article/view/104/103;
http://jodi.ecs.soton.ac.uk/Articles/
v04/i02/Willinsky.

Deniman:2004:MMC

[132] Dave Deniman, Tamara Sumner, Lynne Davis, Sonal Bhushan, and Jackson Fox. Merging metadata and content-based retrieval. Journal of Digital Information, 4(3):??, November 2004. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/105;   http://journals.tdl.
org/jodi/article/view/105/104;
http://jodi.ecs.soton.ac.uk/Articles/
v04/i03/Deniman.

Lempel:2004:CSD

[133] Ronny Lempel, Einat Amitay, David Carmel, Adam Darlow, and Aya Soffer. The connectivity sonar: Detecting site functionality by structural patterns. Journal of Digital Information, 4(3):??, November 2004. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/108;   http://journals.tdl.
org/jodi/article/view/108/107;
http://jodi.ecs.soton.ac.uk/Articles/
v04/i03/Lempel.

Reimer:2004:ICAa

[134] Yolanda Jacobs Reimer and Sarah A. Douglas. Implementation challenges associated with developing a Webbased E-notebook — addendum on related work. Journal of Digital Information, 4(3):??, 2004. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/107;   http://journals.tdl.
org/jodi/article/view/107/106.

Reimer:2004:ICAb

[135] Yolanda Jacobs Reimer and Sarah A. Douglas. Implementation challenges associated with developing a Web-based E-notebook. Journal of Digital Information, 4(3):??, 2004. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/106;   http://journals.tdl.
org/jodi/article/view/106/105.

Binding:2004:KYS

[136] C. Binding and D. Tudhope. KOS at your service: Programmatic access to Knowledge Organisation Systems. Journal of Digital Information, 4(4):??, February 2004. CODEN ???? ISSN

1368-7506. URL http://jodi.tamu.
edu/Articles/v04/i04/Binding.

Smith:2004:BST

[137] T. Smith and M. Zeng. Building semantic tools for concept-based learning spaces. Journal of Digital Information, 4(4):??, January 2004. CODEN ???? ISSN 1368-7506. URL http://jodi.

tamu.edu/Articles/v04/i04/Smith.

Soergel:2004:RTN

[138] D. Soergel, B. Lauser, A. Liang, and F. Fisseha. Reengineering thesauri for new applications: The AGROVOC example. Journal of Digital Information, 4 (4):??, March 2004. CODEN ???? ISSN

1368-7506. URL http://jodi.tamu.
edu/Articles/v04/i04/Soergel.

Tudhope:2004:ENA

[139] D. Tudhope and T. Koch. Editorial: New applications of Knowledge Organization Systems: introduction to a special issue. Journal of Digital Information, 4(4):??, February 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v04/i04/editorial.

Veltman:2004:TSW

[140] K. Veltman. Towards a Semantic Web for culture. Journal of Digital Information, 4(4):??, March 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v04/i04/Veltman.

Vizine-Goetz:2004:VMT

[141] D. Vizine-Goetz, C. Hickey, A. Houghton, and R. Thompson. Vocabulary mapping for terminology services. Journal of Digital Information, 4(4):??, March 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v04/i04/Vizine-Goetz.

Ashman:2004:EFV

[142] H. Ashman and A. Moore. Editorial: Future visions of common-use hypertext: introduction to a special issue. Journal of Digital Information, 5(1):??, July 2004. CODEN ???? ISSN 1368-

                                        7506. URL http://jodi.tamu.edu/
                                        Articles/v05/i01/editorial.

DeBra:2004:NBT

[143] P. De Bra, L. Aroyo, and V. Chepegin. The next big thing: Adaptive Web-based systems. Journal of Digital Information, 5(1):??, May 2004. CODEN ???? ISSN 1368-

                                        7506. URL http://jodi.tamu.edu/
                                        Articles/v05/i01/DeBra.

Brown:2004:IRW

[144] P. Brown and H. Brown. Integrating reading and writing of documents. Journal of Digital Information, 5(1):??, February 2004. CODEN ???? ISSN

                                        1368-7506. URL http://jodi.tamu.
                                        edu/Articles/v05/i01/Brown.

Catanio:2004:UMW

[145] J. Catanio, N. Nnadi, L. Zhang, M. Bieber, and R. Galnares. Ubiquitous metainformation and the WYWWYWI principle. Journal of Digital Information, 5(1):??, April 2004. CODEN ???? ISSN 1368-7506. URL http:

                                        //jodi.tamu.edu/Articles/v05/i01/
                                        Catanio.

Edmonds:2004:PIK

[146] K. Edmonds, J. Blustein, and D. Turnbull. A personal information and knowledge infrastructure integrator. Journal of Digital Information, 5(1):??, May 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v05/i01/Edmonds.

Harper:2004:PWL

[147] S. Harper, C. Goble, and S. Pettitt.

proXimity: Walking the link. Journal of Digital Information, 5(1):??, April 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v05/i01/Harper.

DiIorio:2004:WW

[148] A. Di Iorio and F. Vitali. Writing the Web. Journal of Digital Information, 5 (1):??, May 2004. CODEN ???? ISSN

1368-7506. URL http://jodi.tamu.
edu/Articles/v05/i01/DiIorio.

Mayans:2004:FMT

[149] R. Mayans. The future of mathematical text: A proposal for a new Internet hypertext for mathematics. Journal of Digital Information, 5(1):??, May 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v05/i01/Mayans.

Moore:2004:UHB

[150] A. Moore and T. Brailsford. Unified hyperstructures for bioinformatics: Escaping the application prison. Journal of Digital Information, 5(1):??, May 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v05/i01/Moore.

mschraefel:2004:YGH

[151] m. schraefel, L. Carr, D. De Roure, and W. Hall. You’ve got hypertext. Journal of Digital Information, 5(1):??, July 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v05/i01/schraefel.

Murray-Rust:2004:NBT

[152] P. Murray-Rust and H. Rzepa. The next big thing: From hypermedia to datuments. Journal of Digital Information, 5 (1):??, March 2004. CODEN ???? ISSN

                                        1368-7506. URL http://jodi.tamu.
                                        edu/Articles/v05/i01/Murray-Rust.

Nelson:2004:CDC

[153] T. Nelson. A cosmology for a different computer universe: Data model, mechanisms, virtual machine and visualization infrastructure. Journal of Digital Information, 5(1):??, July 2004. CODEN ???? ISSN 1368-

                                        7506. URL http://jodi.tamu.edu/
                                        Articles/v05/i01/Nelson.

Obendorf:2004:IAP

[154] H. Obendorf. The indirect authoring paradigm — bringing hypertext into the Web. Journal of Digital Information, 5 (1):??, April 2004. CODEN ???? ISSN

                                        1368-7506. URL http://jodi.tamu.
                                        edu/Articles/v05/i01/Obendorf.

Calvo:2004:MCA

[155] R. Calvo, J. Lee, and X. Li. Managing content with automatic document classification. Journal of Digital Information, 5(2):??, June 2004. CODEN ???? ISSN 1368-7506. URL http://jodi.

                                        tamu.edu/Articles/v05/i02/Calvo.

Kong:2004:CFM

[156] X. Kong, L. Liu, and D. Lowe. Critical feature method — A lightweight Web maintenance methodology for SMEs. Journal of Digital Information, 5(2):??, August 2004. CODEN ???? ISSN

                                        1368-7506. URL http://jodi.tamu.
                                        edu/Articles/v05/i02/Kong.

Lowe:2004:EID

[157] D. Lowe. Editorial: Information design models and processes: introduction to a special issue. Journal of Digital Information, 5(2):??, August 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v05/i02/editorial.

Montero:2004:ADT

[158] S. Montero, P. Díaz, and I. Aedo. AriadneTool: A design toolkit for hypermedia applications. Journal of Digital Information, 5(2):??, June 2004. CO-DEN ???? ISSN 1368-7506. URL http:

//jodi.tamu.edu/Articles/v05/i02/
Montero.

Navarro:2004:PAF

[159] A. Navarro, B. Fernández-Manjón, A. Fernández-Valmayor, and J. Sierra. The PlumbingXJ approach for fast prototyping of Web applications. Journal of Digital Information, 5(2):??, June 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v05/i02/Navarro.

Tongrungrojana:2004:WWM

[160] R. Tongrungrojana and D. Lowe.

WIED: A Web modelling language for modelling architectural-level information flows. Journal of Digital Information, 5(2):??, August 2004. CODEN ???? ISSN 1368-7506. URL http:

//jodi.tamu.edu/Articles/v05/i02/
Tongrungrojana.

Bartolo:2004:MID

[161] L. Bartolo, C. Lowe, L. Feng, and B. Patten. MatDL: Integrating digital libraries into scientific practice. Journal of Digital Information, 5(3):??, August 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v05/i03/Bartolo.

Bishop:2004:SCI

[162] A. Bishop, B. Bruce, K. Lunsford, M. Jones, M. Nazarova, D. Linderman, M. Won, P. Heidorn, R. Ramprakash, and A. Brock. Supporting community inquiry with digital resources. Journal of Digital Information, 5(3):??, August 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v05/i03/Bishop.

Coleman:2004:EDL

[163] A. Coleman and T. Sumner. Editorial: Digital libraries and user needs: Negotiating the future. Journal of Digital Information, 5(3):??, September 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v05/i03/editorial.

Foulke:2004:PPC

[164] K. Foulke, N. Milnor, M. Watterworth, and T. Wilsted. The power of partnering: The cooperative creation of digital collections. Journal of Digital Information, 5(3):??, September 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v05/i03/Foulke.

Khoo:2004:UVA

[165] M. Khoo. The use of visual artifacts in the user-centred design of educational digital libraries. Journal of Digital Information, 5(3):??, September 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v05/i03/Khoo.

McMartin:2004:CSM

[166] F. McMartin. Case study: MERLOT:

A model of user involvement in digital library design and implementation. Journal of Digital Information, 5(3):??, September 2004. CODEN ???? ISSN

1368-7506. URL http://jodi.tamu.
edu/Articles/v05/i03/McMartin.

Shreeves:2004:EEU

[167] S. Shreeves and C. Kirkham. Experiences of educators using a portal of aggregated metadata. Journal of Digital Information, 5(3):??, September 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v05/i03/Shreeves.

Ekbia:2004:HIM

[168] H. Ekbia and R. Kling. How IT mediates organizations: Enron and the California Energy Crisis. Journal of Digital Information, 5(4):??, October 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v05/i04/Ekbia.

Lamb:2004:ESA

[169] R. Lamb and S. Johnson. Editorial: Social aspects of digital information in perspective: introduction to a special issue. Journal of Digital Information, 5(4):??, December 2004. CODEN ???? ISSN

1368-7506. URL http://jodi.tamu.
edu/Articles/v05/i04/editorial.

McLean:2004:HOO

[170] P. McLean and S. Johnson. How Oke– Ogun crosses the digital divide — study of a Nigerian rural development project. Journal of Digital Information, 5(4):??, October 2004. CODEN ???? ISSN

1368-7506. URL http://jodi.tamu.
edu/Articles/v05/i04/McLean.

Raisinghani:2004:AIC

[171] M. Raisinghani, A. Benoit, J. Ding, M. Gomez, K. Gupta, V. Gusila, D. Power, and O. Schmedding. Ambient intelligence: Changing forms of humancomputer interaction and their influence on people and organizational environments. Journal of Digital Information, 5 (4):??, August 2004. CODEN ???? ISSN

1368-7506. URL http://jodi.tamu.
edu/Articles/v05/i04/Raisinghani.

Stam:2004:ERI

[172] K. Stam, J. Stanton, and I. Guzman. Employee resistance to information technology change in a social service agency: A membership category approach. Journal of Digital Information, 5(4):??, October 2004. CODEN ???? ISSN 1368-7506. URL http://jodi.

tamu.edu/Articles/v05/i04/Stam.

Esser:2004:FTF

[173] W. Esser. Fault-tolerant fulltext search for large multilingual scientific text corpora. Journal of Digital Information, 6(1):??, October 2004. CODEN ???? ISSN 1368-7506. URL http://jodi.

tamu.edu/Articles/v06/i01/Esser.

Jones:2004:DAL

[174] R. Jones. Designing adaptable learning resources with learning object patterns. Journal of Digital Information, 6(1):??, December 2004. CODEN ???? ISSN 1368-7506. URL http://jodi.

tamu.edu/Articles/v06/i01/Jones.

Poulos:2004:TCT

[175] M. Poulos, S. Papavlasopoulos, and V. Chrissikopoulos. A text categorization technique based on a numerical conversion of a symbolic expression and an onion layers algorithm. Journal of Digital Information, 6(1):??, October 2004. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v06/i01/Poulos.

Jones:2004:SBD

[176] S. Jones, M. Jones, M. Barr, and T. Keegan. Searching and browsing in a digital library of historical maps and newspapers. Journal of Digital Information, 6(2):??, December 2004. CODEN ???? ISSN 1368-7506. URL http://jodi.

tamu.edu/Articles/v06/i02/Jones1.

Klashner:2005:IDE

[177] R. Klashner. ICT and the deregulation of the electric power industry. Journal of Digital Information, 5(4):??, March 2005. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v05/i04/Klashner.

Hemminger:2005:CCV

[178] B. Hemminger, G. Bolas, and D. Schiff.

Capturing content for virtual museums: from pieces to exhibits. Journal of Digital Information, 6(1):??, March 2005. CODEN ???? ISSN 1368-

7506. URL http://jodi.tamu.edu/
Articles/v06/i01/Hemminger.

Hillesund:2005:DTC

[179] T. Hillesund. Digital text cycles:

From medieval manuscripts to modern markup. Journal of Digital Information, 6(1):??, March 2005. CODEN ???? ISSN 1368-7506. URL http:

//jodi.tamu.edu/Articles/v06/i01/
Hillesund.

Bodner:2006:FAE

[180] Richard Bodner and Mark Chignell. A framework for automated evaluation of hypertext search interfaces. Journal of Digital Information, 7(1):??, 2006. CO-DEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/176;   http://journals.tdl.
                                        org/jodi/article/view/176/156.

Francisco-Revilla:2006:CMM

[181] Luis Francisco-Revilla and Frank Shipman. Conflict management in multimodel adaptive hypermedia. Journal of Digital Information, 7(1):??, 2006. CO-DEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/178;   http://journals.tdl.
                                        org/jodi/article/view/178/158.

McRoy:2006:MCA

[182] Susan W. McRoy, Songsak Channarukul, and Syed S. Ali. Multimodal content adaptations for heterogeneous devices. Journal of Digital Information, 7(1):??, 2006. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/177;   http://journals.tdl.
                                        org/jodi/article/view/177/157.

Muntean:2006:EUQ

[183] Cristina Hava Muntean and Jennifer McManis. End-user quality of experience oriented adaptive E-learning system. Journal of Digital Information, 7(1):??, 2006. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/179;   http://journals.tdl.
                                        org/jodi/article/view/179/159.

Mourkoussis:2006:FIA

[184] Nicholaos Mourkoussis, Manjula Patel, and Martin White. A framework for the implementation of application profiles in XML schemas. Journal of Digital Information, 7(2):??, 2006. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/169;   http://journals.tdl.
org/jodi/article/view/169/161.

Pan:2006:BGC

[185] Bing Pan, Helene Hembrooke, Geri Gay, and Gerald C. Gonsalves. Bridging the gap: A conceptual model of the access of digital libraries. Journal of Digital Information, 7(2):??, 2006. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/184;   http://journals.tdl.
org/jodi/article/view/184/162.

Ubois:2006:FMB

[186] Jeff Ubois. Finding Murphy Brown:

How accessible are historic television broadcasts? Journal of Digital Information, 7(2):??, 2006. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/172;   http://journals.tdl.
org/jodi/article/view/172/155.

Kong:2007:WST

[187] Xiaoying Kong, Li Liu, and David Lowe.

A Web system trace model and its application to Web design. Journal of Digital Information, 8(1):??, 2007. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/202;   http://journals.tdl.
org/jodi/article/view/202/164.

Ambacher:2007:GAD

[188] Bruce I. Ambacher. Government archives and the digital repository audit checklist. Journal of Digital Information, 8(2):??, 2007. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/190;   http://journals.tdl.
org/jodi/article/view/190/171.

Chavez:2007:SMR

[189] Robert Chavez, Gregory Crane, Anne Sauer, Alison Babeu, Adrian Packel, and Gabriel Weaver. Services make the repository. Journal of Digital Information, 8(2):??, 2007. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/195;   http://journals.tdl.
org/jodi/article/view/195/179.

Crabtree:2007:BRP

[190] Jonathan David Crabtree and Darrell Donakowski. Building relationships project update 2007. Journal of Digital Information, 8(2):??, 2007. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/191;   http://journals.tdl.
org/jodi/article/view/191/172.

Dobratz:2007:NCC

[191] Susanne Dobratz, Astrid Schoger, and Stefan Strathmann. The nestor catalogue of criteria for trusted digital repository evaluation and certification. Journal of Digital Information, 8(2):??, 2007. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/199;   http://journals.tdl.
org/jodi/article/view/199/180.

Efron:2007:MUO

[192] Miles Efron. Metadata use in OAI-compliant institutional repositories. Journal of Digital Information, 8(2):??, 2007. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/196;   http://journals.tdl.
org/jodi/article/view/196/169.

Kim:2007:MIF

[193] Jihyun Kim. Motivating and impeding factors affecting faculty contribution to institutional repositories. Journal of Digital Information, 8(2):??, 2007. CO-DEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/193;   http://journals.tdl.
org/jodi/article/view/193/177.

Lee:2007:DCT

[194] Christopher A. Lee and Helen R. Tibbo.

Digital curation and trusted repositories: Steps toward success. Journal of Digital Information, 8(2):??, 2007. CODEN ???? ISSN 1368-

7506. URL http://journals.tdl.
org/jodi/article/view/229/183.

Markey:2007:NCI

[195] Karen Markey, Beth St. Jean, Soo Young Rieh, Elizabeth Yakel, Jihyun Kim, and Yong-Mi Kim. Nationwide census of institutional repositories: Preliminary findings. Journal of Digital Information, 8(2):??, 2007. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/194;   http://journals.tdl.
org/jodi/article/view/194/170.

Moore:2007:AVT

[196] Reagan W. Moore and MacKenzie Smith. Automated validation of trusted digital repository assessment criteria. Journal of Digital Information, 8(2):??, 2007. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/198;   http://journals.tdl.
org/jodi/article/view/198/181.

Thibodeau:2007:IYB

[197] Kenneth Thibodeau. If you build it, will it fly? criteria for success in a digital repository. Journal of Digital Information, 8(2):??, 2007. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/197;   http://journals.tdl.
org/jodi/article/view/197/174.

Branny:2007:ASE

[198] Emilia Branny. Automatic summary evaluation based on text grammars. Journal of Digital Information, 8(3):??, 2007. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/232;   http://journals.tdl.
org/jodi/article/view/232/186.

Cristea:2007:AHS

[199] Alexandra Cristea, Craig Stewart, Tim Brailsford, and Paul Cristea. Adaptive hypermedia system interoperability: a ‘real world’ evaluation. Journal of Digital Information, 8(3):??, 2007. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/235;   http://journals.tdl.
org/jodi/article/view/235/192.

Cristea:2007:TGA

[200] Alexandra Cristea, David Smits, and Paul de Bra. Towards a generic adaptive hypermedia platform: a conversion case study. Journal of Digital Information, 8(3):??, 2007. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/231;   http://journals.tdl.
org/jodi/article/view/231/184.

Duarte:2007:DAD

[201] Carlos Duarte and Luís Carriço. Developing an adaptive digital talking book player with FAME. Journal of Digital Information, 8(3):??, 2007. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/236;   http://journals.tdl.
org/jodi/article/view/236/193.

Kolbitsch:2007:DAC

[202] Josef Kolbitsch, Christian Safran, and Hermann Maurer. Dynamic adaptation of content and structure in electronic encyclopaedias. Journal of Digital Information, 8(3):??, 2007. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/237;   http://journals.tdl.
org/jodi/article/view/237/191.

Maneewatthana:2007:API

[203] Thanyalak Maneewatthana, Gary Wills, and Wendy Hall. Adaptive personal information environment based on the Semantic Web. Journal of Digital Information, 8(3):??, 2007. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/233;   http://journals.tdl.
org/jodi/article/view/233/187.

Rykowski:2007:VWS

[204] Jarogniew Rykowski. Virtual Web services. Journal of Digital Information, 8(3):??, 2007. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/234;   http://journals.tdl.
                                        org/jodi/article/view/234/188.

Wilde:2007:PSD

[205] Erik Wilde. Personalization of shared data: The ShaRef approach. Journal of Digital Information, 8(3):??, 2007. CO-DEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/241;   http://journals.tdl.
                                        org/jodi/article/view/241/194.

Bradley:2008:PMD

[206] John Douglas Bradley. Pliny: A model for digital support of scholarship. Journal of Digital Information, 9(1):??, 2008. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/209;   http://journals.tdl.
                                        org/jodi/article/view/209/198.

Indratmo:2008:ROS

[207] J. Indratmo and Julita Vassileva. A review of organizational structures of personal information management. Journal of Digital Information, 9(1):??, 2008. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/251;   http://journals.tdl.
                                        org/jodi/article/view/251/200.

Kennedy:2008:NQG

[208] Marie R. Kennedy. Nine questions to guide you in choosing a metadata schema. Journal of Digital Information, 9(1):??, 2008. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/226;   http://journals.tdl.
org/jodi/article/view/226/205.

Verbert:2008:AFT

[209] Katrien Verbert, Xavier Ochoa, and Erik Duval. The ALOCOM framework: Towards scalable content reuse. Journal of Digital Information, 9(1):??, 2008. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/225;   http://journals.tdl.
org/jodi/article/view/225/199.

Cornish:2008:DSN

[210] Alan Kevin Cornish and Trevor James Bond. Developing and sustaining the northwest digital archives. Journal of Digital Information, 9(2):??, 2008. CO-DEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/245;   http://journals.tdl.
org/jodi/article/view/245/201.

Heckner:2008:TTA

[211] Markus Heckner, Susanne Mühlbacher, and Christian Wolff. Tagging tagging. analysing user keywords in scientific bibliography management systems. Journal of Digital Information, 9(2):??, 2008. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/246;   http://journals.tdl.
org/jodi/article/view/246/208.

Pomerantz:2008:DLS

[212] Jeffrey Pomerantz. Digital (library services) and (digital library) services. Journal of Digital Information, 9(2):??, 2008. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/227;   http://journals.tdl.
                                        org/jodi/article/view/227/210.

Delort:2009:ACS

[213] Jean-Yves Delort. Automatically characterizing salience using readers’ feedback. Journal of Digital Information, 10(1):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/268;   http://journals.tdl.
                                        org/jodi/article/view/268/274.

Goh:2009:ISI

[214] Dion Goh. Introduction to the special issue of the journal of digital information digital libraries and user-generated content. Journal of Digital Information, 10 (1):??, 2009. CODEN ???? ISSN 1368-

                                        7506. URL http://journals.tdl.
                                        org/jodi/article/view/450/285.

Lee:2009:TSI

[215] Chei Sian Lee, Dion Hoe-Lian Goh, Khasfariyati Razikin, and Alton Y. K. Chua. Tagging, sharing and the influence of personal experience. Journal of Digital Information, 10(1):??, 2009. CO-DEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/275;   http://journals.tdl.
                                        org/jodi/article/view/275/275.

Maslov:2009:CCW

[216] Alexey Maslov, Adam Mikeal, Katherine Weimer, and John Leggett. Cooperation or control? Web 2.0 and the digital library. Journal of Digital Information, 10(1):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/271;   http://journals.tdl.
org/jodi/article/view/271/276.

Trant:2009:SST

[217] J. Trant. Studying social tagging and folksonomy: A review and framework. Journal of Digital Information, 10(1):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/269;   http://journals.tdl.
org/jodi/article/view/269/278.

Trant:2009:TFA

[218] J. Trant. Tagging, folksonomy and art museums: Early experiments and ongoing research. Journal of Digital Information, 10(1):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/270;   http://journals.tdl.
org/jodi/article/view/270/277.

Drachsler:2009:IGU

[219] Hendrik Drachsler, Hans G. K. Hummel, and Rob Koper. Identifying the goal, user model and conditions of recommender systems for formal and informal learning. Journal of Digital Information, 10(2):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/442;   http://journals.tdl.
org/jodi/article/view/442/279.

Dron:2009:LSS

[220] Jon Dron and Terry Anderson. Lost in social space: Information retrieval issues in Web 1.5. Journal of Digital Information, 10(2):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/443;   http://journals.tdl.
                                        org/jodi/article/view/443/280.

Duval:2009:SIS

[221] Erik Duval, Riina Vuorikari, and Nikos Manouselis. Special issue on social information retrieval for technology enhanced learning. Journal of Digital Information, 10(2):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/444/281.

Klerkx:2009:VSB

[222] Joris Klerkx and Erik Duval. Visualising social bookmarks. Journal of Digital Information, 10(2):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/445;   http://journals.tdl.
                                        org/jodi/article/view/445/282.

Tang:2009:PVP

[223] Tiffany Y. Tang and Gordon McCalla. The pedagogical value of papers: a collaborative-filtering based paper recommender. Journal of Digital Information, 10(2):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/446;   http://journals.tdl.
                                        org/jodi/article/view/446/283.

Vuorikari:2009:EAM

[224] Riina Vuorikari and Xavier Ochoa. Exploratory analysis of the main characteristics of tags and tagging of educational resources in a multi-lingual context. Journal of Digital Information, 10(2):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/447;   http://journals.tdl.
org/jodi/article/view/447/284.

Bernstein:2009:SCH

[225] Mark Bernstein. Shadows in the cave:

hypertext transformations. Journal of Digital Information, 10(3):??, 2009. CO-DEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/714;   http://journals.tdl.
org/jodi/article/view/714/488.

Hsieh:2009:SVP

[226] Haowei Hsieh and Frank Shipman. Supporting visual problem solving in spatial hypertext. Journal of Digital Information, 10(3):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/173;   http://journals.tdl.
org/jodi/article/view/173/486.

Kolb:2009:OSS

[227] David Kolb. Other spaces for spatial hypertext. Journal of Digital Information, 10(3):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/171;   http://journals.tdl.
org/jodi/article/view/171/485.

Meintanis:2009:EPI

[228] Konstantinos A. Meintanis and Frank M.

Shipman. Expressing personal interpretations of music collections in spatial hypertext. Journal of Digital Information, 10(3):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/702;   http://journals.tdl.
org/jodi/article/view/702/480.

Rosenberg:2009:HOA

[229] Jim Rosenberg. Hypertext in the open air: A systemless approach to spatial hypertext. Journal of Digital Information, 10(3):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/175;   http://journals.tdl.
                                        org/jodi/article/view/175/487.

Shipman:2009:ISI

[230] Frank Shipman and Jim Rosenberg. Introduction to special issue on spatial hypertext. Journal of Digital Information, 10(3):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/715/489.

Garrido:2009:CTM

[231] Piedad Garrido, Jesus Tramullas, and Manuel Coll. CONDITOR1: Topic maps and DITA labelling tool for textual documents with historical information. Journal of Digital Information, 10(4):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/264;   http://journals.tdl.
                                        org/jodi/article/view/264/202.

Perugini:2009:ETI

[232] Saverio Perugini, Mary E. Pinney, Naren Ramakrishnan, Manuel A. Pérez-Quiñones, and Mary Beth Rosson. Exploring out-of-turn interactions with Websites. Journal of Digital Information, 10(4):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/205;   http://journals.tdl.
                                        org/jodi/article/view/205/206;

http://journals.tdl.org/jodi/article/
view/205/528.

Roche:2009:TWM

[233] Mathieu Roche and Yves Kodratoff.

Text and Web mining approaches in order to build specialized ontologies. Journal of Digital Information, 10(4):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/645;   http://journals.tdl.
org/jodi/article/view/645/506.

Wolfram:2009:LSD

[234] Dietmar Wolfram and Iris Xie. A longitudinal study of database usage within a general audience digital library. Journal of Digital Information, 10(4):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/304;   http://journals.tdl.
org/jodi/article/view/304/505.

Barreau:2009:SIS

[235] Deborah Barreau. Special issue on the social and psychological aspects of personal information management. Journal of Digital Information, 10(5):??, 2009. CODEN ???? ISSN 1368-

7506. URL http://journals.tdl.
org/jodi/article/view/748/527.

Coughlan:2009:DPI

[236] Tim Coughlan and Peter Johnson.

Designing personal information management systems for creative practitioners. Journal of Digital Information, 10(5):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/

                                        view/440;   http://journals.tdl.
                                        org/jodi/article/view/440/540.

Elsweiler:2009:URB

[237] David Elsweiler, Mark Baillie, and Ian Ruthven. On understanding the relationship between recollection and refinding. Journal of Digital Information, 10(5):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/436;   http://journals.tdl.
                                        org/jodi/article/view/436/542.

Hardof-Jaffe:2009:SOS

[238] Sharon Hardof-Jaffe, Arnon Hershkovitz, Hama Abu-Kishk, Ofer Bergman, and Rafi Nachmias. Students’ organization strategies of personal information space. Journal of Digital Information, 10(5):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/438;   http://journals.tdl.
                                        org/jodi/article/view/438/541.

Baruzzo:2009:TSD

[239] Andrea Baruzzo, Paolo Casoto, Prasad Challapalli, Antonina Dattolo, Nirmala Pudota, and Carlo Tasso. Toward semantic digital libraries: Exploiting Web 2.0 and semantic services in cultural heritage. Journal of Digital Information, 10(6):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/688;   http://journals.tdl.
                                        org/jodi/article/view/688/576.

Byrne:2009:PHC

[240] Kate Byrne. Putting hybrid cultural data on the Semantic Web. Journal of Digital Information, 10(6):??, 2009. CO-DEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/700;   http://journals.tdl.
org/jodi/article/view/700/579.

Clough:2009:EDS

[241] Paul D. Clough, Neil Ireson, and Jennifer Marlow. Extending domain-specific resources to enable semantic access to cultural heritage data. Journal of Digital Information, 10(6):??, 2009. CO-DEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/698;   http://journals.tdl.
org/jodi/article/view/698/578.

Koutsomitropoulos:2009:MSD

[242] Dimitrios A. Koutsomitropoulos, Georgia D. Solomou, and Theodore S. Papatheodorou. Metadata and semantics in digital object collections: A case-study on CIDOC-CRM and Dublin Core and a prototype implementation. Journal of Digital Information, 10(6):??, 2009. CO-DEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/693;   http://journals.tdl.
org/jodi/article/view/693/577.

Larson:2009:SII

[243] Martha Larson, Kate Fernie, and Johan Oomen. Special issue on information access to cultural heritage. Journal of Digital Information, 10(6):??, 2009. CO-DEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/803;   http://journals.tdl.
org/jodi/article/view/803/580.

Lunn:2009:UNT

[244] Brian Kirkegaard Lunn. User needs in television archive access: Acquiring knowledge necessary for system design. Journal of Digital Information, 10(6):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/685;   http://journals.tdl.
                                        org/jodi/article/view/685/560.

Marx:2009:AIA

[245] Maarten Marx. Advanced information access to parliamentary debates. Journal of Digital Information, 10(6):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/668;   http://journals.tdl.
                                        org/jodi/article/view/668/552.

Ordelman:2009:TAD

[246] Roeland Ordelman, Willemijn Heeren, Marijn Huijbregts, Franciska de Jong, and Djoerd Hiemstra. Towards affordable disclosure of spoken heritage archives. Journal of Digital Information, 10(6):??, 2009. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/687;   http://journals.tdl.
                                        org/jodi/article/view/687/562.

Aschenbrenner:2010:TOR

[247] Andreas Aschenbrenner, Tobias Blanke, Marc W. Küster, and Wolfgang Pempe. Towards an open repository environment. Journal of Digital Information, 11(1):??, 2010. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/758;   http://journals.tdl.
                                        org/jodi/article/view/758/646.

Downs:2010:SAL

[248] Robert R. Downs and Robert S.

Chen. Self-assessment of a longterm archive for interdisciplinary scientific data as a trustworthy digital repository. Journal of Digital Information, 11(1):??, 2010. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/753;   http://journals.tdl.
org/jodi/article/view/753/642.

Gerber:2010:AEV

[249] Anna Gerber and Jane Hunter. Authoring, editing and visualizing compound objects for literary scholarship. Journal of Digital Information, 11(1):??, 2010. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/755;   http://journals.tdl.
org/jodi/article/view/755/643.

Kelly:2010:ASC

[250] Julia Kelly and Louise Letnes. AgEcon search: A case study on the differences between operating a subject repository and an institutional repository. Journal of Digital Information, 11(1):??, 2010. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/733;   http://journals.tdl.
org/jodi/article/view/733/639.

Maslov:2010:AOR

[251] Alexey Maslov, James Creel, Adam Mikeal, Scott Phillips, John Leggett, and Mark McFarland. Adding OAI– ORE support to repository platforms. Journal of Digital Information, 11(1):??, 2010. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/749;   http://journals.tdl.
                                        org/jodi/article/view/749/640.

McDonald:2010:RTR

[252] Robert H. McDonald and Tyler O. Walters. Restoring trust relationships within the framework of collaborative digital preservation federations. Journal of Digital Information, 11(1):??, 2010. CO-DEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/757;   http://journals.tdl.
                                        org/jodi/article/view/757/645.

Rajasekar:2010:PBD

[253] Arcot Rajasekar, Reagan Moore, Mike Wan, and Wayne Schroeder. Policybased distributed data management systems. Journal of Digital Information, 11(1):??, 2010. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/756;   http://journals.tdl.
                                        org/jodi/article/view/756/644.

Sefton:2010:ITE

[254] Peter Sefton and Jim Downing. ICEtheorem — end to end semantically aware eResearch infrastructure for theses. Journal of Digital Information, 11(1):??, 2010. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/754;   http://journals.tdl.
                                        org/jodi/article/view/754/647;
                                        http://journals.tdl.org/jodi/article/
                                        view/754/657.

Warner:2010:AIS

[255] Simeon Warner. Author identifiers in scholarly repositories. Journal of Digital Information, 11(1):??, 2010. CO-DEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/752;   http://journals.tdl.
org/jodi/article/view/752/641.

Davies:2011:MSW

[256] Stephen Davies, Chris Donaher, Jesse Hatfield, and Jessica Zeitz. Making the Semantic Web usable: interface principles to empower the layperson. Journal of Digital Information, 12(1):??, 2011. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/746;   http://journals.tdl.
org/jodi/article/view/746/1182.

Breure:2011:RIP

[257] Leen Breure, Hans Voorbij, and Maarten Hoogerwerf. Rich Internet publications: “show what you tell”. Journal of Digital Information, 12(1), 2011. CO-DEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/1606;   http://journals.tdl.
org/jodi/article/view/1606/1738.

Knutov:2011:GAF

[258] Evgeny Knutov, Paul De Bra, and Mykola Pechenizkiy. Generic adaptation framework: a process-oriented perspective. Journal of Digital Information, 12(1), 2011. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/965;   http://journals.tdl.
org/jodi/article/view/965/1739.

ORiordan:2011:EOW

[259] Adrian O’Riordan and Aliver O’Mahony.

Engineering an Open Web Syndication Interchange with discovery and recommender capabilities. Journal of Digital Information, 12(1), 2011. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/962;   http://journals.tdl.
org/jodi/article/view/962/1744.

Abrams:2011:CMS

[260] Stephen Abrams, Patricia Cruse, John Kunze, and David Minor. Curation micro-services: A pipeline metaphor for repositories. Journal of Digital Information, 12(2):??, 2011. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/1605;   http://journals.tdl.
org/jodi/article/view/1605/1766.

Losch:2011:BDA

[261] Mathias Lösch, Ulli Waltinger, Wolfram Horstmann, and Alexander Mehler. Building a DDC-annotated corpus from OAI metadata. Journal of Digital Information, 12(2), 2011. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/1765;   http://journals.tdl.
org/jodi/article/view/1765/1767.

Pickton:2011:PRC

[262] Miggie Pickton, Debra Morris, Stephanie Meece, Simon Coles, and Steve Hitchcock. Preserving repository content: practical tools for repository managers. Journal of Digital Information, 12(2), 2011. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/1767;   http://journals.tdl.
org/jodi/article/view/1767/1768.

Kaplan:2011:ADO

[263] Deborah Kaplan, Anne Sauer, and Eliot Wilczek. Archival description in OAI–ORE. Journal of Digital Information, 12(2), 2011. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/1814;   http://journals.tdl.
org/jodi/article/view/1814/1769.

Aschenbrenner:2011:DAR

[264] Andreas Aschenbrenner, Harry Enke, Thomas Fischer, and Jens Ludwig. Diversity and interoperability of repositories in a grid curation environment. Journal of Digital Information, 12(2), 2011. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/1896;   http://journals.tdl.
org/jodi/article/view/1896/1770.

Doorenbosch:2011:IRL

[265] Paul Doorenbosch and Barbara Sierman. Institutional repositories, long term preservation and the changing nature of scholarly publications. Journal of Digital Information, 12(2), 2011. CO-DEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/1764;   http://journals.tdl.
org/jodi/article/view/1764/1851.

Sefton:2011:RPE

[266] Peter Sefton and Duncan Dickinson.

Repositories post 2010: embracing heterogeneity in AWE, the Academic Working Environment. Journal of Digital Information, 12(2), 2011. CODEN ???? ISSN 1368-7506. URL http:

//journals.tdl.org/jodi/article/
view/1766;   http://journals.tdl.
org/jodi/article/view/1766/2165;

                                         http://journals.tdl.org/jodi/article/
                                        view/1766/2166;   http://journals.
                                        tdl.org/jodi/article/view/1766/
                                        2167.

Covey:2011:RCI

[267] Denise Troll Covey. Recruiting content for the institutional repository: The barriers exceed the benefits. Journal of Digital Information, 12(3):??, 2011. CO-DEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/2068;   http://journals.tdl.
                                        org/jodi/article/view/2068/1757.

Gartner:2011:ISC

[268] Richard Gartner. Intermediary schemas for complex XML applications: an example from research information management. Journal of Digital Information, 12(3), 2011. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/2069;   http://journals.tdl.
                                        org/jodi/article/view/2069/2086.

Yahiaoui:2011:RCM

[269] Leila Yahiaoui, Yannick Prié, Zizette Boufaida, and Pierre-Antoine Champin. Redocumenting computer-mediated activity from its traces: a modelbased approach for narrative construction. Journal of Digital Information, 12(3), 2011. CODEN ???? ISSN 1368-7506. URL http:

                                        //journals.tdl.org/jodi/article/
                                        view/2088;   http://journals.tdl.
                                        org/jodi/article/view/2088/1756.

Bell:2011:DWR

[270] Graeme Baxter Bell. Digital whistleblowing in restricted environments. Journal of Digital Information, 12 (3), 2011. CODEN ???? ISSN

1368-7506. URL http://journals.
tdl.org/jodi/article/view/2136;
http://journals.tdl.org/jodi/article/
view/2136/2087.

Mikhail:2012:DMI

[271] Youssef Mikhail, Noha Adly, and Magdy Nagi. DAR: a modern institutional repository with a scalability twist. Journal of Digital Information, 13(1):??, 2012. CODEN ???? ISSN 1368-7506.

Tzoc:2012:DVN

[272] Elias Tzoc. Document viewers for non-Born-digital files in DSpace. Journal of Digital Information, 13(1), 2012. CO-DEN ???? ISSN 1368-7506.

Lewis:2012:RSR

[273] Stuart Lewis, Kim Shepherd, Yin Yin Latt, Andrea Schweer, and Adam Field. Repository as a Service (RaaS). Journal of Digital Information, 13(1), 2012. CODEN ???? ISSN 1368-7506.

Adams:2012:CWI

[274] Sam Adams and Peter Murray-Rust.

Chempound — a Web 2.0-inspired repository for physical science data. Journal of Digital Information, 13(1), 2012. CODEN ???? ISSN 1368-7506.

Hswe:2012:BCC

[275] Patricia Hswe, Michael J. Giarlo, Michelle Belden, Kevin Clair, Daniel Coughlin, and Linda Klimczyk. Building a community of curatorial practice at Penn State: a case study. Journal of Digital Information, 13(1), 2012. CO-DEN ???? ISSN 1368-7506.

Brewer:2012:RDP

[276] James Brewer, Tracy Popp, and Joy Perrin. REDDNET and digital preservation in the Open Cloud: Research at Texas Tech University Libraries on long-term archival storage. Journal of Digital Information, 13(1), 2012. CODEN ???? ISSN 1368-7506.

Waddington:2012:CMR

[277] Simon Waddington, Richard Green, and Chris Awre. CLIF: Moving repositories upstream in the content lifecycle. Journal of Digital Information, 13(1), 2012. CODEN ???? ISSN 1368-7506.

Waddington:2012:KRS

[278] Simon Waddington, Jun Zhang, Gareth Knight, Mark Hedges, Jens Jensen, and Roger Downing. Kindura: Repository services for researchers based on hybrid clouds. Journal of Digital Information, 13(1), 2012. CODEN ???? ISSN 1368-7506.

Olendorf:2012:BLH

[279] Robert Olendorf and Steve Koch. Beyond the low hanging fruit: Data services and archiving at the University of New Mexico. Journal of Digital Information, 13(1), 2012. CODEN ???? ISSN 1368-7506.

Awre:2012:BHT

[280] Chris Awre and Tom Cramer. Building the Hydra together: Enhancing repository provision through multi-institution collaboration. Journal of Digital Information, 13(1), 2012. CODEN ???? ISSN 1368-7506.

Nuernberg:2012:CIT

[281] Peter Nuernberg, John Leggett, and Mark McFarland. Cloud as infrastructure at the Texas Digital Library. Journal of Digital Information, 13(1), 2012. CODEN ???? ISSN 1368-7506.

Hedges:2012:SCE

[282] Mark Hedges, Tobias Blanke, Stella Fabiane, Gareth Knight, and Eric Liao. Sheer curation of experiments: Data, process, provenance. Journal of Digital Information, 13(1), 2012. CODEN ???? ISSN 1368-7506.

Hedges:2012:FED

[283] Mark Hedges, Mike Haft, and Gareth Knight. FISHNet: encouraging data sharing and reuse in the freshwater science community. Journal of Digital Information, 13(1), 2012. CODEN ???? ISSN 1368-7506.

Bertazzo:2012:PDA

[284] Matteo Bertazzo and Angela Di Iorio. Preserving and delivering audiovisual content integrating Fedora Commons and MediaMosa. Journal of Digital Information, 13(1), 2012. CODEN ???? ISSN 1368-7506.

Ball:2012:VRD

[285] Alexander Ball, Mansur Darlington, Thomas Howard, Chris McMahon, and Steve Culley. Visualizing research data records for their better management. Journal of Digital Information, 13(1), 2012. CODEN ???? ISSN 1368-7506.